### v1

# A shared-revenue Bertrand game

Abstract.
We introduce and analyze a variation of the Bertrand game in which the revenue is shared between two players. This game models situations in which one economic agent can provide goods/services to consumers either directly or through an independent seller/contractor in return for a share of the revenue. We analyze the equilibria of this game, and show how they can predict different business outcomes as a function of the players’ costs and the transferred revenue shares. Importantly, we identify game parameters for which independent sellers can simultaneously increase the original player’s payoff while increasing consumer surplus. We then extend the shared-revenue Bertrand game by considering the shared revenue proportion as an action and giving the independent seller an outside option to sell elsewhere. This work constitutes a first step towards a general theory for how partnership and sharing of resources between economic agents can lead to more efficient markets and improve the outcomes of both agents as well as consumers.
conference: ; ;
1. Introduction
1.1. Motivation
Many modern businesses partner with independent service providers to increase efficiency. One method through which this is done is for a company to let a separate company provide services/goods to customers in exchange for a share of the revenue. The independent service provider in turn may benefit from the wider visibility it gains through access to the original company’s pool of customers. This is seen in a number of real-world scenarios, including but not limited to:
  1. (A)
    Consider a large store which allows independent sellers to use some of its shelf space to sell products in return for a portion of the selling price. The large store may of course obtain the same products from a vendor and sell them on its own, and the independent seller may set up its own store, but depending on the costs for each party, both parties might benefit from having the independent seller’s products available in the larger store.
  2. (B)
    Consider a construction and maintenance company, with a large pool of customers, providing a variety of services such as plumbing, repairs, painting services, etc. The company can either provide a specific service on its own, or refer the customer to a contractor who sets the price for the service, and pays the original company a referral fee.
  3. (C)
    Consider a ridesharing or taxi platform. When a passenger requests a ride, they have the option of transporting the passenger with either (i) a self-driving car, owned by the ridesharing platform, or (ii) an independent human driver, whom the ridesharing platform charges a fee for each passenger they refer.
  4. (D)
    A large airline could either operate a short regional flight with their own aircraft or allow a regional partner to operate the flight.
  5. (E)
    A healthcare provider can either provide specialist services in-house or refer their patients to an external contractor, charging a referral fee.
1.2. Overview
In this work we provide a game theoretic model for reasoning about the interaction between agents in scenarios such as these with a variation of the Bertrand pricing game (Bertrand, 1883). For concreteness, we will focus our intuition and analysis on scenario (A) from Section 1.1, though our model can explain phenomena in domains far beyond retail sales. In accordance with scenario (A), the “shared-revenue Bertrand game” is played between a retailer, and an independent goods/services provider which we will refer to as the independent seller (seller). Both players sell an identical good/service, with each agent having a different cost. Under the revenue sharing program, whenever the independent seller fulfills demand, a portion of the revenue, called the referral fee, is transferred from the independent seller to the retailer. In Section 2, we set up the game in more detail, then in Section 3 we identify the Nash equilibria of this game. In Section 4, we interpret and analyze the equilibria of the shared-revenue Bertrand game in depth with different costs and fee structures.
The referral fee is a constant parameter of the shared-revenue Bertrand game of Section 2. However, to make the game more realistic, in Section 5 we introduce the “fee optimization game,” which models the referral fee as an action for the retailer that is chosen before playing the shared-revenue Bertrand game. Finally, in Section 6, we introduce the “outside option game”: after observing the referral fee chosen by the retailer, the independent seller can choose to sell on their own instead of through the retailer.
The equilibria of the “fee optimization game” reveal that the retailer always prefers to allow the independent seller to fulfill demand in equilibrium, setting the referral fee strategically so as to maximize their payoff. There is a sweet spot for the optimal fee such that it is low enough to stimulate economic activity while high enough to maximize payoff. In the fee optimization game of Section 5, we still observe some unrealistic equilibria in which the retailer sets such a high fee that the independent seller must fulfill demand at their effective cost (thus achieving no payoff), which leads us to consider the “outside option game” in Section 6. If the independent seller’s threat of selling on their own is credible, the unrealistic equilibria of Section 5 will no longer be observed, as the retailer must set the referral fee lower to ensure the independent seller doesn’t exercise their outside option. Importantly, our analysis demonstrates that the retailer will frequently be incentivized to set a fee under which both the independent seller increases its payoff compared to selling independently, and the price to the customer is decreased relative to the price either the retailer or the independent seller would have set independently. Our results therefore suggest that such a revenue sharing mechanism can not only increase social welfare, but also simultaneously improve outcomes for both players as well as consumers.
In each of the models, we examine if the revenue sharing program increases the retailer’s payoff compared to simply fulfilling demand on their own, and how it affects the end cost to the customer.
1.3. Related work
Our model is an extension of the standard Bertrand game (Bertrand, 1883) in which two sellers of an identical product with the same cost set their prices independently. As we justify in Appendix B, we consider the regime in which the sellers’ costs are asymmetric, which is a natural extension of the standard Bertrand game, studied for instance in (Blume, 2003; Kartik, 2011; Demuynck, 2019).
One field in which revenue sharing programs as described in Section 1.1 have been studied is in the independent contractor literature. Much of this literature analyzes relationships with independent contractors through the framework of principal agent theory (Coats, 2002; Chang, 2014). The revenue-sharing program can also be seen as a form of “make-or-buy” decision for the retailer (Coase, 1937; Williamson, 1975; Tadelis, 2002). Slightly less obviously, another field in which revenue sharing programs have been studied is optimal taxation theory, where the referral fee can be viewed as the retailer imposing a tax on the independent seller. Viewed in this light, our results about optimal referral fees resonate strongly with optimal taxation theories such as (Laffer, 2004; Stern, 1976; Stiglitz, 1987).
Finally, the idea of an outside option that we introduce in Section 6 is common in negotiation and bargaining theory (Binmore et al., 1989; Osborne and Rubinstein, 1990), which makes sense because the participation of the independent seller in the retailer’s revenue-sharing program can be seen as a negotiation of a contract. In fact, there is even an “outside option principle” (Watson, 2020), which informally states than an outside option must be credible in order to have an effect on the Nash equilibria of the game; our results are consistent with this.
2. Preliminaries
Our model is a game in which two players – the retailer ($r$), and the independent seller ($s$), sell an identical product. The actions in the game are the price each player chooses to charge, and we assume the players move simultaneously. Formally, we define the following game:
Shared-Revenue Bertrand Game
Parameters
Retailer’s cost $c_{r}$, seller’s cost $c_{s}$, referral fee $\alpha$
Gameplay
  1. (1)
    The retailer and seller simultaneously choose prices, yielding a price configuration $(p_{r},p_{s})$
The payoffs for the players are
(1) $\bar{\pi}_{r}$ $=(\bar{p}_{r}-\bar{c}_{r})\bar{q}_{r}+\alpha\bar{p}_{s}\bar{q}_{s}$
(2) $\bar{\pi}_{s}$ $=(\bar{p}_{s}-\bar{c}_{s}-\alpha\bar{p}_{s})\bar{q}_{s}$
where $\bar{p}$, $\bar{c}$, $\bar{q}$ and $\alpha$ denote prices, costs, quantities sold and the referral fee, respectively, and subscripts denote the players. This game can be viewed as a modification of the Bertrand game, where the players have different costs, and a portion of the independent seller’s revenue is shared with the retailer as part of the service provided by the retailer.
To allow for a concrete mathematical analysis we assume demand is linear in price, and that the agent with the lower price fulfills all demand. When both players set the same price, we assume they split the market where the retailer fulfills $\beta\in[0,1]$ of the demand and the independent seller fulfills the rest. This assumption is made for mathematical completeness, although in Proposition 1 we show that no split market equilibrium exists, making the choice of $\beta$ irrelevant. Concretely, the retailer’s demand curve is given by
(3) $\bar{q}_{r}(\bar{p}_{r},\bar{p}_{s})=\begin{cases}Q_{0}-K^{(q)}\bar{p}_{r}&\bar{p}_{r}<\bar{p}_{s}\\ \beta(Q_{0}-K^{(q)}\bar{p}_{r})&\bar{p}_{r}=\bar{p}_{s}\\ 0&\bar{p}_{r}>\bar{p}_{s}\end{cases}$
where $Q_{0}$ and $K^{(q)}$ are constant parameters. The demand curve for the independent seller is analogous, but $\beta$ is replaced by $(1-\beta)$.
To reduce the number of parameters in the model we will use the following normalization, and throughout the rest of the paper present the results in terms of unitless variables
(4) $p=\bar{p}\frac{K^{(q)}}{Q_{0}};\quad c=\bar{c}\frac{K^{(q)}}{Q_{0}};\quad q=\bar{q}\frac{1}{Q_{0}};\quad\pi=\bar{\pi}\frac{K^{(q)}}{Q_{0}^{2}}.$
Under this normalization, the demand at price $p=0$ is 1, and the demand at price $p=1$ is 00. Equations (1) and (2) remain unchanged if we remove bars, and (3) reduces to
(5) $q_{r}(p_{r},p_{s})=\begin{cases}1-p_{r}&p_{r}<p_{s}\\ \beta(1-p_{r})&p_{r}=p_{s}\\ 0&p_{r}>p_{s}\end{cases}$
2.1. Necessary assumptions
In order for the game to be nontrivial, we must impose the following constraints:
  1. (I)
    There exists a price at which both the retailer and independent seller could achieve a nonzero payoff when fulfilling demand ($10<c_{s},c_{r}<$ 1).
  2. (II)
    The seller’s cost is less than the retailer’s cost ($c_{s}<c_{r}$).
  3. (III)
    The referral fee takes some, but not all of the independent seller’s revenue ($\alpha\in(0,1)$).
In Appendix B, we formally show that the game degenerates if these are violated, but the importance of each condition should be fairly intuitive. Overall, the revenue-sharing program should be thought of as for products that the retailer could conceivably sell themselves, but the independent seller is able to do so more efficiently.
2.2. Important prices and payoffs
We now introduce several important prices and corresponding payoffs which will be used throughout this paper. We also include a full notation table in Appendix A, and include derivations of the key prices and fees in Appendix D.
As we will later prove in Proposition 1, we only need to focus on cases where either the retailer or the independent seller fulfill all demand, and can ignore any split market scenario. When we write a payoff $\pi$ with two subscripts, the first denotes the the player whose payoff we refer to, and the second denotes the player who fulfills all demand. Thus, $\pi_{r,r}(p_{r})$ and $\pi_{s,r}(p_{r})$ are the retailer’s and the independent seller’s payoffs when the retailer fulfills all the demand, respectively. Similarly $\pi_{r,s}(p_{s})$ and $\pi_{s,s}(p_{s})$ are the retailer’s and the independent seller’s payoffs when the independent seller fulfills all the demand, respectively. Explicitly, it follows from (1), (2), and (5) that
(6) $\pi_{r,r}(p_{r})=(p_{r}-c_{r})(1-p_{r})\ ;$ $\pi_{s,r}(p_{r})=0$
(7) $\pi_{r,s}(p_{s})=\alpha p_{s}(1-p_{s})\ ;$ $\pi_{s,s}(p_{s})=((1-\alpha)p_{s}-c_{s})(1-p_{s})$
In Figure 1 we plot the payoffs of the players as a function of price for two choices of game parameters $(c_{r},c_{s},\alpha)$. Solid and dashed lines denote the players’ payoff when the retailer or independent seller fulfill all demand, respectively, while blue and red lines denote the payoffs to the retailer and the independent seller, respectively.
Figure 1. Typical payoff curves and important price points
Figure 1 also illustrates five important price points and their corresponding payoffs which will be used throughout this paper. We denote by $p_{r^{*}}$ and $p_{s^{*}}$ the prices each player would set to maximize their own payoff if they fulfilled all demand, and call these the retailer’s or independent seller’s optimal selling price. Differentiating the players’ single agent payoffs $\pi_{r,r}$ and $\pi_{s,s}$ and equating to zero in (36) and (40) gives the values of these prices and their corresponding payoffs –
(8) $p_{r^{*}}\equiv\frac{1+c_{r}}{2};$ $\pi_{r,r}(p_{r^{*}})=\frac{(1-c_{r})^{2}}{4}$
(9) $p_{s^{*}}\equiv\frac{1+\frac{c_{s}}{1-\alpha}}{2};$ $\pi_{s,s}(p_{s^{*}})=\begin{cases}\frac{1-\alpha}{4}\left(1-\frac{c_{s}}{1-\alpha}\right)^{2}&\alpha<1-c_{s}\\ 0&\alpha\geq 1-c_{s}\end{cases}$
We denote by $p_{r_{\text{ind}}}$ and $p_{s_{\text{ind}}}$ the prices at which the retailer and the independent seller are indifferent to which player fulfills demand, i.e. $\pi_{r,r}(p_{r_{\text{ind}}})=\pi_{r,s}(p_{r_{\text{ind}}})$ and $\pi_{s,r}(p_{s_{\text{ind}}})=\pi_{s,s}(p_{s_{\text{ind}}})$. Solving these equations in (38) and (42) yields the following prices and payoffs –
(10) $p_{r_{\text{ind}}}\equiv\frac{c_{r}}{1-\alpha};$ $\pi_{r,r}(p_{r_{\text{ind}}})=\begin{cases}\frac{\alpha c_{r}}{1-\alpha}\left(1-\frac{c_{r}}{1-\alpha}\right)&\alpha<1-c_{r}\\ 0&\alpha\geq 1-c_{r}\end{cases}$
(11) $p_{s_{\text{ind}}}\equiv\frac{c_{s}}{1-\alpha};$ $\pi_{s,s}(p_{s_{\text{ind}}})=0$
We first note that $\pi_{s,s}(p_{s_{\text{ind}}})=0$, and therefore the independent seller’s indifference price is also its breakeven price.
More interestingly, while the retailer’s breakeven price is still its cost, $c_{r}$, its indifference price is scaled by $(1-\alpha)^{-1}$, but at that price its payoff is nonzero. This is the main feature which differentiates the shared-revenue Bertrand game from the classical Bertrand game, as the retailer can choose to fulfill demand on its own or allow the independent seller to fulfill demand, making revenue from the fee. As a result, this game gives rise to a more complex equilibrium structure which depends on the fees and costs of the players. Note also that because we assume $c_{s}<c_{r}$ (Section 2.1), we have $p_{s_{\text{ind}}}<p_{r_{\text{ind}}}$.
Next, we introduce the price at which the retailer’s payoff when the independent seller fulfills demand is equal to its optimal payoff had it fulfilled demand on its own, i.e.
$\pi_{r,r}(p_{r^{*}})=\pi_{r,s}(p^{\dagger})$
(12) $p^{\dagger}\equiv\frac{1+\sqrt{1-\frac{(1-c_{r})^{2}}{\alpha}}}{2}$
$p^{\dagger}$ is the solution of a quadratic equation which we derive in (58), and we define it as the larger of the two possible solutions. Furthermore, $p^{\dagger}$ is only well-defined when $\alpha\geq(1-c_{r})^{2}$, and therefore we only show it in the right subplot of Figure 1.
3. Equilibrium analysis
3.1. Market split
For simplicity, we will only consider pure strategies in this paper; thus, a strategy profile can be characterized fully by a tuple $(p_{r},p_{s})\in\mathcal{S}_{p}\equiv(0,1)\times(0,1)$ denoting the prices chosen by the retailer and independent seller. We first prove, in Proposition 1, that in all equilibria, one player takes the entire market.
Proposition 0.
Let $(p_{r},p_{s})=(p,p)$ be an equilibrium price configuration, with the equilibrium quantities sold by each player $q_{r}=\beta(1-p)$ and $q_{s}=(1-\beta)(1-p)$. Then $\beta\in\{0,1\}$.
Proof sketch.
To prove the Proposition we show that when the market is split, the retailer’s payoff is a weighted average of $\pi_{r,r}$ and $\pi_{r,s}$, and one of them is larger than the split market payoff. Thus, the retailer can always increase their payoff by either lowering or raising the price by an infinitesimal amount, fulfilling demand themselves or allowing the independent seller to do so respectively. See Appendix C for the complete proof. ∎
While this result is not surprising given that in nearly all standard non-symmetric game theory models, in equilibrium only one player fulfills all demand, this result is clearly not expected to hold in the real world. Game theory results may mimic the world more closely and arrive at more realistic equilibria in which players split demand by modeling factors such as incomplete information and brand loyalty, but these are beyond the scope of this work.
3.2. Nash equilibrium derivation
In light of Proposition 1, we can denote all equilibria in the game by either unequal prices $(p_{r},p_{s})\in\mathcal{S}_{p}$ or a shared price $(p,x)\in(0,1)\times\{r,s\}$, where $x$ denotes the player who fulfills the demand. In this subsection, we aim to provide an intuitive sketch of the Nash equilibria.
To narrow our focus, let’s consider shared-price equilibria $(p,s)$ where the independent seller fulfills demand at price $p$ for now. In order for $(p,s)$ to be an Nash equilibrium, the following conditions must be satisfied:
  1. (1)
    0 = $\pi_{s,r}(p)\leq\pi_{s,s}(p)$
  2. (2)
    $\sup_{p^{\prime}\in[0,p)}\left[\pi_{s,s}(p^{\prime})\right]\leq\pi_{s,s}(p)$
  3. (3)
    $\sup_{p^{\prime}\in[0,p)}\left[\pi_{r,r}(p^{\prime})\right]\leq\pi_{r,s}(p)$
If condition 1 doesn’t hold, the independent seller can achieve a higher payoff by increasing its price and letting the retailer fulfill demand. If condition 2 doesn’t hold, the independent seller can increase its payoff by lowering their price and continuing to fulfill all demand. If condition 3 doesn’t hold, then there is some price lower than $p$ the retailer can set to take the entire market and increase its payoff.
Figure 2. Shared-price Nash equilibria where the independent seller fulfills demand
In Figure 2, we highlight in yellow the Nash equilibrium prices where the independent seller fulfills demand. In the left subplot of Figure 2, the Nash equilibria occur at prices $p\in[p_{s_{\text{ind}}},p_{r_{\text{ind}}}]$, which satisfies all of the Nash equilibrium conditions:
  1. (1)
    The seller achieves nonnegative payoff because $p\geq p_{s_{\text{ind}}}$.
  2. (2)
    The independent seller’s payoff would decrease if they set a price $p^{\prime}<p$ because $p\leq p_{s^{*}}$.
  3. (3)
    The retailer would not prefer to fulfill demand themselves because $p\leq p_{r_{\text{ind}}}$.
In the right subplot of Figure 2, the Nash equilibria occur at prices $p\in[p_{s_{\text{ind}}},p^{\dagger}]$. Just as in the left subplot, conditions 1 and 2 are satisfied because $p\geq p_{s_{\text{ind}}}$ and $p\leq p_{s^{*}}$. Condition 3 is satisfied because $p\leq p^{\dagger}$. Notice that for the payoff curves in the right subplot, this is a stronger condition than $p\leq p_{r_{\text{ind}}}$. This stronger condition is necessary because if $p>p^{\dagger}$, the retailer would prefer to set price $p_{r^{*}}<p$ and achieve a payoff of
$\pi_{r,r}(p_{r^{*}})=\pi_{r,s}(p^{\dagger})>\pi_{r,s}(p)$
In Appendix D.2, we carry forth this intuition and translate conditions 1-3 into feasible intervals of prices for any $\alpha\in(0,1)$ rather than just for the specific referral fees in Figure 2. For $(p,s)$ to be a Nash equilibrium, we find that the price must satisfy
(13) $p\in\left[p_{s_{\text{ind}}},\min\{p_{r_{\text{ind}}},p_{r^{*}},p_{s^{*}}\}\right]\cup\left[\max\{p_{r^{*}},p_{s_{\text{ind}}}\},\min\{p^{\dagger},p_{s^{*}}\}\right]$
We acknowledge that it’s still not clear what the equilibrium interval of prices look like for each possible choice of $\alpha$, namely what lower and upper bounds are attained for each possible referral fee – this warrants a longer discussion which we defer to Section 3.3. For instance, $p^{\dagger}$ is not well-defined for small $\alpha$, but we will soon see in (20) that whenever $\left[\max\{p_{r^{*}},p_{s_{\text{ind}}}\},\min\{p^{\dagger},p_{s^{*}}\}\right]$ is nonempty, $\alpha$ is high enough for $p^{\dagger}$ to be well-defined. For now, simply keep in mind that (13) is sufficient only for shared-price Nash equilibria where the independent seller fulfills demand.
In Appendix D.1, we derive the conditions analogous to (13) for shared-price equilibria $(p,r)$ where the retailer fulfills demand and find the following:
(14) $p\in\left[p_{r_{\text{ind}}},\min\{p_{s_{\text{ind}}},p_{r^{*}},p_{s^{*}}\}\right]$
Upon further inspection of (14), we notice a subtle problem. Recall that $p_{s_{\text{ind}}}<p_{r_{\text{ind}}}$ always because $c_{s}<c_{r}$. Thus, the interval $\left[p_{r_{\text{ind}}},\min\{p_{s_{\text{ind}}},p_{r^{*}},p_{s^{*}}\}\right]$ is empty for all $\alpha$, which implies there are no shared price-equilibria where the retailer fulfills demand!
This completes the first step of our analysis for shared-price equilibria, but we also need to consider unequal-price equilibria. Indeed, we consider equilibria where $p_{r}<p_{s}$ in Appendix D.3 and equilibria where $p_{s}<p_{r}$ in Appendix D.4. Interested readers can find conditions similar to (13) and (14) in the corresponding appendices; we also include a summary of all Nash equilibria with both equal and unequal prices in Section 3.4.
3.3. Important fees for transitions between Nash equilibria
In this section, we answer the aforementioned question of what the interval in (13) looks like for different values of $\alpha$. To do this, we will introduce five threshold fees that delineate the transitions between when different Nash equilibria can occur (skip ahead to Figure 3 for a preview). Similar analysis as in this section for unequal price equilibria can be found in Appendices D.3 and D.4.
Consider first the left interval of (13), $p\in\left[p_{s_{\text{ind}}},\min\{p_{r_{\text{ind}}},p_{r^{*}},p_{s^{*}}\}\right]$. We know that $p_{s_{\text{ind}}}<p_{r_{\text{ind}}}$ always. We have $p_{s_{\text{ind}}}\leq p_{s^{*}}=\frac{1}{2}(1+p_{s_{\text{ind}}})$ so long as $p_{s_{\text{ind}}}\leq 1$, which occurs when $\alpha\leq 1-c_{s}$. The condition for $p_{s_{\text{ind}}}\leq p_{r^{*}}$ is not quite as clear, so we introduce the retailer optimal price feasibility fee $\alpha_{r^{*}}$ (derived in (46)) –
(15) $p_{s_{\text{ind}}}\leq p_{r^{*}}\iff\alpha\leq\alpha_{r^{*}}\ ;$ where $\text{where }\alpha_{r^{*}}\equiv 1-\frac{2c_{s}}{1+c_{r}}$
The name comes from the fact that $p_{r^{*}}\leq p_{s_{\text{ind}}}$ is necessary for any equilibrium where the retailer fulfills demand at their optimal price, which we prove in Appendix D.3. Note also that $\alpha_{r^{*}}\leq 1-c_{s}$, which we show in (52). Thus, the left interval of (13) is nonempty if and only if $\alpha\leq\alpha_{r^{*}}$.
Now, we turn our attention to the value of $\min\{p_{r_{\text{ind}}},p_{r^{*}},p_{s^{*}}\}$, the upper bound of the left interval of (13). This motivates the optimality switching fee $\alpha_{s^{*},r^{*}}$ (derived in (52)), the seller optimal price feasibility fee $\alpha_{s^{*}}$ (derived in (50)), and the retailer $p^{\dagger}$ relevance fee $\alpha^{\dagger}_{r}$ (derived in (49)) –
(16) $p_{s^{*}}\leq p_{r^{*}}\iff\alpha\leq\alpha_{s^{*},r^{*}}\ ;$ where $\text{where }\alpha_{s^{*},r^{*}}\equiv\frac{c_{r}-c_{s}}{c_{r}}$
(17) $p_{r_{\text{ind}}}\leq p_{s^{*}}\iff\alpha\leq\alpha_{s^{*}}\ ;$ where $\text{where }\alpha_{s^{*}}\equiv 1-2c_{r}+c_{s}$
(18) $p_{r_{\text{ind}}}\leq p_{r^{*}}\iff\alpha\leq\alpha^{\dagger}_{r}\ ;$ where $\text{where }\alpha^{\dagger}_{r}\equiv\frac{1-c_{r}}{1+c_{r}}$
Analogously to $\alpha_{r^{*}}$, we will soon find that $\alpha\geq\alpha_{s^{*}}$ is necessary for an equilibrium where the independent seller fulfills demand at $p_{s^{*}}$. Intersecting the inequalities in (15), (16), (17), and (18) is sufficient to determine what the left interval of (13) looks like in each region of parameter space. We will omit the algebra here and instead opt to summarize this in Section 3.4. However, throughout the course of the derivation in Appendix D, an important condition emerges that determines the relative ordering of $\alpha_{s^{*},r^{*}}$, $\alpha_{s^{*}}$, and $\alpha^{\dagger}_{r}$. We show in (54) and (55) that this seller optimal feasibility cost $c_{s^{*}}$ satisfies the following –
(19) $\alpha_{s^{*}}\leq\alpha^{\dagger}_{r}\iff\alpha^{\dagger}_{r}\leq\alpha_{s^{*},r^{*}}\iff c_{s}\leq c_{s^{*}}\ ;$ where $\text{where }c_{s^{*}}\equiv\frac{2c_{r}^{2}}{1+c_{r}}$
Finally, let’s turn our attention to the right interval of (13), $p\in\left[\max\{p_{r^{*}},p_{s_{\text{ind}}}\},\min\{p^{\dagger},p_{s^{*}}\}\right]$. Fortunately, we have already determined the relative ordering of each of the prices in the right interval, except for $p^{\dagger}$. For these conditions involving $p^{\dagger}$, we introduce the $p^{\dagger}$ relevance fees for the retailer and seller (derived in (59) and (62)):
(20) $p^{\dagger}\leq p_{r^{*}}\iff\alpha\leq\alpha^{\dagger}_{r}\ ;$ where $\text{where }\alpha^{\dagger}_{r}\equiv\frac{1-c_{r}}{1+c_{r}}$
(21) $p_{s_{\text{ind}}}\leq p^{\dagger}\iff\alpha\leq\alpha^{\dagger}_{s}\ ;$
Note that we already introduced $\alpha^{\dagger}_{r}$ in (18). The notation arises because we will show that $\alpha^{\dagger}_{r}\leq\alpha\leq\alpha^{\dagger}_{s}$ is necessary for any equilibrium where demand is fulfilled at $p^{\dagger}$. We don’t include a closed form here for $\alpha^{\dagger}_{s}$ because it is uninsightful, but we remark that $\alpha^{\dagger}_{s}\in[\alpha_{r^{*}},1-c_{s}]$ and that a closed form is in (62). Finally, the relative ordering of $p_{s^{*}}$ and $p^{\dagger}$ is complex, and there unfortunately is no special $\hat{\alpha}$ for which $p_{s^{*}}\leq p^{\dagger}\iff\alpha\leq\hat{\alpha}$. Thus, we will leave that condition unsimplified for now.
3.4. Summary of Nash equilibria
Considering all possible values of $\alpha$ and all possible price configurations, we summarize the Nash equilibria of the game below. We denote sets of equilibrium price configurations as $(S_{r},S_{s})\subset\mathcal{S}_{p}$ and singletons without the set notation, specifically $x\equiv\{x\}$. Recall there are no shared-price equilibria where the retailer fulfills demand, thus all sets of shared-price equilibria $(S_{r},S_{s})=(p_{s},[\cdot,\cdot])$ implicitly assume that the independent seller is fulfilling demand at price $p_{s}\in[\cdot,\cdot]$, not the retailer.
Equilibrium Price Configurations of Shared-Revenue Bertrand Game
Notation
We denote Nash equilibrium strategies by price configurations $(p_{r},p_{s})\in(S_{r},S_{s})\subset\mathcal{S}_{p}$.
Case (i), $c_{s}\leq c_{s^{*}}$:
(22) $(p_{r},p_{s})\in\begin{cases}(p_{s},\left[p_{s_{\text{ind}}},p_{r_{\text{ind}}}\right])&\alpha\leq\alpha_{s^{*}}\\ (\left[p_{s^{*}},1\right],p_{s^{*}})\cup(p_{s},\left[p_{s_{\text{ind}}},p_{s^{*}}\right])&\alpha_{s^{*}}\leq\alpha\leq\alpha_{s^{*},r^{*}}\\ (\left[p_{s^{*}},1\right],p_{s^{*}})\cup(p_{s},\left[p_{s_{\text{ind}}},p_{s^{*}}\right])&\left[\alpha_{s^{*},r^{*}}\leq\alpha\leq\alpha^{\dagger}_{s}\right]\land\left[p_{s^{*}}\leq p^{\dagger}\right]\\ (p_{s},\left[p_{s_{\text{ind}}},p^{\dagger}\right])&\left[\alpha_{s^{*},r^{*}}\leq\alpha\leq\alpha^{\dagger}_{s}\right]\land\left[p^{\dagger}\leq p_{s^{*}}\right]\\ (p_{r^{*}},\left[p^{\dagger},1\right])&\alpha_{r^{*}}\leq\alpha\end{cases}$
Case (ii), $c_{s}\geq c_{s^{*}}$:
(23) $(p_{r},p_{s})\in\begin{cases}(p_{s},\left[p_{s_{\text{ind}}},p_{r_{\text{ind}}}\right])&\alpha\leq\alpha^{\dagger}_{r}\\ (\left[p_{s^{*}},1\right],p_{s^{*}})\cup(p_{s},\left[p_{s_{\text{ind}}},p_{s^{*}}\right])&\left[\alpha^{\dagger}_{r}\leq\alpha\leq\alpha^{\dagger}_{s}\right]\land\left[p_{s^{*}}\leq p^{\dagger}\right]\\ (p_{s},\left[p_{s_{\text{ind}}},p^{\dagger}\right])&\left[\alpha^{\dagger}_{r}\leq\alpha\leq\alpha^{\dagger}_{s}\right]\land\left[p^{\dagger}\leq p_{s^{*}}\right]\\ (p_{r^{*}},\left[p^{\dagger},1\right])&\alpha_{r^{*}}\leq\alpha\end{cases}$
4. Refining and interpreting equilibria
In (22) and (23) of the last section, we found a proliferation of Nash equilibrium price configurations in $\mathcal{S}_{p}$ that could be observed given the costs and the value of $\alpha$. However, to help us interpret the Nash equilibria, we would like to discern which price configurations are the “most likely” to occur.
To this end, we introduce some criteria to distinguish Nash equilibria:
  1. (1)
    Admissibility - A Nash equilibrium in which any player plays a weakly dominated strategy [^2] Player $x$’s strategy $s_{x}$ is weakly dominated by strategy $s_{x}^{\prime}$ if (a) for all strategies of other players $s_{y}$, $\pi_{x}(s_{x}^{\prime},s_{y})\geq\pi_{x}(s_{x},s_{y})$, and (b) there exists a strategy of the other players $s_{y}$ such that $\pi_{x}(s_{x}^{\prime},s_{y})>\pi_{x}(s_{x},s_{y})$ (Rasmusen, 1989). is inadmissible (Govindan and Wilson, 2016).
  2. (2)
    Relative Pareto optimality - We prefer Nash equilibria that are Pareto optimal [^3] Nash equilibrium $X$ is Pareto suboptimal relative to Nash equilibrium $Y$ if (a) all players weakly prefer, and (b) at least one player strictly prefers $X$ over $Y$ (where “prefer” means achieving a higher payoff) (Rasmusen, 1989). relative to all other Nash equilibria. [^4] Note that a relatively Pareto optimal Nash equilibrium is not necessarily on the Pareto frontier of the game, because it need not be Pareto optimal in the set of all strategies, only in the set of Nash equilibria.
After these two refinements, we will still be left with some intervals of equilibrium prices. However, most of these are inconsequential; in Section 4.3 we will see that these refinements allow us to identify who will fulfill demand and the price at which they will sell in equilibrium.
4.1. Admissibility
The admissibility criterion allows us to rule out any weakly dominated strategies for the retailer and independent seller. For instance, in the left subplot of Figure 2, all Nash equilibrium prices $p_{r}^{\prime}\in[p_{s_{\text{ind}}},p_{r_{\text{ind}}})$ are weakly dominated by $p_{r_{\text{ind}}}$ for the retailer. Indeed, for all $p_{r}^{\prime}<p_{r_{\text{ind}}}$, we have $\pi_{r,s}(p_{r}^{\prime})>\pi_{r,r}(p_{r}^{\prime})$ by construction of $p_{r_{\text{ind}}}$. Thus, if the seller sets a price $p_{s}\in(p_{r}^{\prime},p_{r_{\text{ind}}})$, then
(24) $\pi_{r}(p_{r}^{\prime},p_{s})=\pi_{r,r}(p_{r}^{\prime})<\pi_{r,r}(p_{r_{\text{ind}}})=\pi_{r,s}(p_{r_{\text{ind}}})<\pi_{r,s}(p_{s})=\pi_{r}(p_{r_{\text{ind}}},p_{s})$
which implies that the retailer would have preferred to set price $p_{r_{\text{ind}}}$ rather than $p_{r}^{\prime}$. Thus, the only admissible Nash equilibrium in the left subplot of Figure 2 is $p=p_{r_{\text{ind}}}$.
In Appendix E.1, we more thoroughly consider the retailer and independent seller’s strategies, and find the strategies that are not weakly dominated are
(25) $p_{r}\in\left[\min\left\{p_{r_{\text{ind}}},p_{r^{*}}\right\},\max\left\{p_{r_{\text{ind}}},p_{r^{*}}\right\}\right];\quad p_{s}\in\left[p_{s_{\text{ind}}},p_{s^{*}}\right]$
The intuition for the independent seller and the retailer is the same – we can eliminate all prices that are not between their indifference and optimal prices – however, for the retailer we do not always have that $p_{r_{\text{ind}}}\leq p_{r^{*}}$. Given the admissible strategies in (25), we can simply intersect the sets of admissible prices with the Nash equilibrium price configurations we found in (22) and (23). For instance, in the right subplot of Figure 2, we find that the admissible Nash equilibria are
$\left[p_{s_{\text{ind}}},p^{\dagger}\right]\cap\left[p_{r^{*}},p_{r_{\text{ind}}}\right]\cap\left[p_{s_{\text{ind}}},p_{s^{*}}\right]=\left[p_{r^{*}},p^{\dagger}\right]$
We perform these intersections throughout all of parameter space in Appendix E.1, and include a complete summary of admissible equilibria in (70), (71), and (72).
4.2. Relative Pareto optimality
The relative Pareto optimality criterion allows us to further eliminate one admissible equilibrium, which occurs in the following case:
$(p_{r},p_{s})\in\left(p_{r^{*}},\left[p^{\dagger},p_{s^{*}}\right]\right)\cup\left(p_{s},\left[p_{s_{\text{ind}}},p^{\dagger}\right]\right)\ ;$ if $\text{if }\alpha_{r^{*}}\leq\alpha\leq\alpha^{\dagger}_{s}\text{ and }p^{\dagger}\leq p_{s^{*}}$
The equilibrium where the retailer fulfills demand at $p_{r^{*}}$ is Pareto suboptimal relative to the shared-price equilibrium $p^{\dagger}$ where the independent seller fulfills demand. Indeed, by construction of $p^{\dagger}$, in both cases the retailer achieves the same payoff of $\pi_{r,r}(p_{r^{*}})=\pi_{r,s}(p^{\dagger})$ so they are indifferent. However, because $\alpha\leq\alpha^{\dagger}_{s}$ implies that $p_{s_{\text{ind}}}\leq p^{\dagger}$, the independent seller achieves a positive payoff when fulfilling demand at $p^{\dagger}$ while they achieve a payoff of 00 when allowing the retailer to fulfill demand at $p_{r^{*}}$. Thus, they strictly prefer fulfilling demand at $p^{\dagger}$ rather than allowing the retailer to fulfill demand at $p_{r^{*}}$.
As a remark, we note that a reasonable argument could be made for preferring the equilibrium where the retailer fulfills demand at $p_{r^{*}}$. In this case $p_{r^{*}}<p^{\dagger}$, so despite the Pareto optimality argument, the customer will be able to buy the product at a lower price if the retailer fulfills demand at $p_{r^{*}}$ instead of the independent seller fulfilling demand at $p^{\dagger}$. With this lowest price criterion, $p_{r^{*}}$ would in fact be the most likely equilibrium in this case, as $p_{r^{*}}<p_{s_{\text{ind}}}$ as well. However, whatever we choose does not substantially affect the results, and the relative Pareto optimality criteria will continue to be sensible for us as we generalize the game, thus we proceed with this assumption.
4.3. Equilibrium outcomes
Even after applying these two refinements, inspecting (70), (71), and (72) we can still see that we are left with many intervals of prices. However, in order to interpret the real-world implications of the equilibria, it is prudent to consider the equilibrium outcomes: at what price the customer will purchase the product and from whom they will buy. Indeed, letting the prices be $p\in P\subset[0,1]$ and the fulfiller of demand be $x\in\{r,s\}$, we find the following equilibria:
Equilibrium Outcomes of Shared-Revenue Bertrand Game
Notation
We denote an equilibrium outcome by $(P,x)\subset(0,1)\times\{r,s\}$, indicating that there is an admissible, Pareto optimal Nash equilibrium where player $x$ fulfills demand at any price $p\in P\subset(0,1)$.
Case (i), $c_{s}\leq c_{s^{*}}$:
(26) $(P,x)=\begin{cases}(p_{r_{\text{ind}}},s)&\alpha\leq\alpha_{s^{*}}\\ (p_{s^{*}},s)&\alpha_{s^{*}}\leq\alpha\leq\alpha_{s^{*},r^{*}}\\ \left(\left[\max\left\{p_{r^{*}},p_{s_{\text{ind}}}\right\},\min\left\{p_{s^{*}},p^{\dagger}\right\}\right],s\right)&\alpha_{s^{*},r^{*}}\leq\alpha\leq\alpha^{\dagger}_{s}\\ (p_{r^{*}},r)&\alpha^{\dagger}_{s}\leq\alpha\end{cases}$
Case (ii), $c_{s}\geq c_{s^{*}}$:
(27) $(P,x)=\begin{cases}(p_{r_{\text{ind}}},s)&\alpha\leq\alpha^{\dagger}_{r}\\ \left(\left[\max\left\{p_{r^{*}},p_{s_{\text{ind}}}\right\},\min\left\{p_{s^{*}},p^{\dagger}\right\}\right],s\right)&\alpha^{\dagger}_{r}\leq\alpha\leq\alpha^{\dagger}_{s}\\ (p_{r^{*}},r)&\alpha^{\dagger}_{s}\leq\alpha\end{cases}$
Viewing the equilibria in the light of (26) and (27) is incredibly helpful for intepretation; the remainder of Section 4 is dedicated to analyzing the outcomes rather than the price configurations.
A quick glance at (26) and (27) shows that when $\alpha$ is sufficiently high, the independent seller is “priced out” of the market and the retailer will fulfill demand at their single agent price as if the independent seller did not exist. When $\alpha$ is small, the independent seller sells at the retailer’s indifference price, illustrating that they have to prioritize preventing the retailer from undercutting them over maximizing their own payoff. However, when $\alpha$ is in a certain sweet spot, the independent seller may be able to sell at their optimal price in equilibrium. We will see soon that everybody wins in this regime – if the retailer gets to choose the referral fee, they would prefer to choose a referral fee in this sweet spot rather than one that is too high or too low.
4.4. Fee and cost structure analysis
Our results in (26) and (27) show that given a point in the parameter space $(c_{r},c_{s},\alpha)$, we can predict which player will fulfill demand and the price(s) at which they will fulfill demand in an admissible, Pareto optimal Nash equilibrium. In this section, we build intuition about how the parameter space is divided into regions based on these possible equilibria and discuss the business implications for the retailer, both in terms of the retailer’s payoff and in terms of the effect on the prices. Importantly, we will contrast these metrics with the alternative of not having the revenue sharing option available to outside sellers – i.e. compare them with the price the retailer would have set and its payoff in the absence of an independent seller.
Figure 3. Equilibrium outcomes as a function of $\alpha$ vs $c_{r}$ (left) and $\alpha$ vs $c_{s}$ (right).
In Figure 3 we plot two cross sections of the parameter space, one plotting $c_{r}$ vs $\alpha$ at a fixed $c_{s}=0.4$ and one with $c_{s}$ vs $\alpha$ at a fixed $c_{r}=0.6$. The different colors correspond to the admissible, Pareto optimal equilibrium for each set of parameters, with the “continuum” corresponding to the interval of prices $\left(p\in\left[\max\left\{p_{r^{*}},p_{s_{\text{ind}}}\right\},\min\left\{p_{s^{*}},p^{\dagger}\right\}\right],s\right)$.
Notice that when $c_{s}\geq c_{s^{*}}$ there is no value of $\alpha$ that yields an equilibrium $(p_{s^{*}},s)$ in the red region – this is where the $c_{s^{*}}$ notation comes from. Additionally, notice that we may have that $\alpha_{s^{*}}<0$, in which case the green region is not feasible as we cannot have $\alpha\leq\alpha_{s^{*}}<0$.
4.4.1. $(p_{r^{*}},r)$ The retailer fulfills demand at its single agent price
When $\alpha>\alpha^{\dagger}_{s}$ (blue region), the fee is too high for the independent seller to fulfill demand, and the game reduces to the retailer “playing” alone and charging its single agent price. This is the least interesting equilibrium if we wish to consider the retailer’s interaction with the independent sellers, but is interesting to be contrasted with all other equilibria as it is the only equilibrium at which the retailer fulfills demand.
4.4.2. A continuum of equilibria in which the independent seller fulfills demand at a price higher than the retailer’s single agent price
When $\alpha\in\left[\max\left\{\alpha_{s^{*},r^{*}},\alpha^{\dagger}_{r}\right\},\alpha^{\dagger}_{s}\right]$, there are a continuum of equilibrium prices at which the independent seller could fulfill demand (gray region). Note that the independent seller’s price is at least $p_{r^{*}}$ in this region of parameter space, thus the price to the end customer increases relative to a world in which the revenue sharing program did not exist. However, in this region, the retailer achieves at least as much payoff as they would achieve fulfilling demand on their own without the revenue sharing program. Intuitively, this equilibrium is “safe” for the retailer because they will never make less than their single agent payoff, but it comes at the expense of potentially charging customers a higher price.
4.4.3. $(p_{r_{\text{ind}}},s)$ The independent seller fulfills demand at the retailer’s indifference price, $p_{r_{\text{ind}}}$
When $\alpha\leq\min\left\{\alpha_{s^{*}},\alpha^{\dagger}_{r}\right\}$, the independent seller fulfills demand at $p=p_{r_{\text{ind}}}$ (green region). In this region there is a trade-off between our two metrics – the retailer’s payoff is lower, but the price for the customer is also lower compared to the retailer selling in a single agent game. Furthermore, the existence of this equilibrium implies that the retailer’s option of selling directly to customer protects the customer from the independent seller charging too high a price when the fee is low.
4.4.4. $(p_{s^{*}},s)$ The independent seller fulfills demand at their single agent price, $p_{s^{*}}$
When $\alpha\in\left[\alpha_{s^{*}},\alpha_{s^{*},r^{*}}\right]$, we have an equilibrium in which the independent seller fulfills demand at its optimal price $p_{s^{*}}$ (red region). In this regime, we always have $p_{s^{*}}\leq p_{r^{*}}$, thus the customers are able to purchase at a lower price than they would be able to without the revenue sharing program. Of course, the independent seller prefers this equilibrium because the fee is fairly low and they are able to sell at their optimal price. Finally, it’s even possible that there exists some $\alpha$ in the red region such that the retailer achieves a payoff higher than their optimal single agent payoff, yielding a win-win-win situation for the customers, independent seller, and retailer.
5. Optimizing the referral fee
In Section 4.4, we identified an equilibrium that is a win-win-win (red region) for the customers, independent seller, and retailer, along with a “safe” equilibrium (gray region) for the retailer to achieve a payoff no less than its single agent payoff. While we have made an intuitive argument about which equilibria are preferable, the next step in our analysis is to generalize the game to allow the retailer to choose the referral fee as an action. With this, we will build a more accurate and formal model to answer the question of which equilibria are preferable to the retailer. Precisely, we study the following sequential game:
Fee Optimization Game
Parameters
Retailer’s cost $c_{r}$, seller’s cost $c_{s}$
Gameplay
  1. (1)
    The retailer chooses the referral fee $\alpha\in(0,1)$
  2. (2)
    The retailer and the independent seller play the shared-revenue Bertrand game with parameters $(\alpha,c_{r},c_{s})$
Examining Figure 3, we can think of fixing a cost on the horizontal axis and taking a slice, for instance drawing a vertical line corresponding to $c_{r}=0.9$ on the left plot. With this, the retailer now gets to choose whatever value of $\alpha$ they’d like, which is a point along the vertical line $c_{r}=0.9$, effectively getting to choose which type of equilibrium they would like to observe. We illustrate such a slice in Figure 4 by coloring the background of the plot according to which equilibrium in Figure 3 is observed. As one might expect, we will find that the retailer prefers the red/gray equilibria to the green/blue equilibria. However, despite potentially being a win-win-win, we will find that the retailer does not always choose the red equilibrium when it exists because they may be able to “win more” in the gray region at the expense of customers and/or the independent seller.
5.1. Refined equilibria of fee optimization game
We will impose the same refinements on the Nash equilibria of the sequential game as we did in Section 4 – namely, that they must be admissible and Pareto optimal relative to other equilibria. Admissibility has no effect on the equilibria of the sequential game, and Pareto optimality yields us the natural refinement that the retailer will choose the minimum $\alpha$ necessary to attain their maximum possible equilibrium payoff. More formal justification can be found in Appendix F.1.
We will impose the further refinement of subgame perfection (Rasmusen, 1989), which is equivalent to backward induction in this game of perfect, complete information. In order to achieve this, the retailer and independent seller must choose a price configuration that is an admissible, Pareto optimal equilibrium of the shared-revenue Bertrand subgame. However, there are many choices of such price configurations. We will thus define an equilibrium strategy profile $\rho$ for the subgame. Formally, we require that $\rho$ is a strategy profile with the property that for every possible $\alpha,c_{r},c_{s}$, $\rho(\alpha,c_{r},c_{s})\in\mathcal{S}_{p}$ is an admissible, Pareto optimal equilibrium price configuration of the shared-revenue Bertrand game as outlined in (70), (71), and (72) [^5] This function $\rho$ is necessary for subgame perfection and could come from anywhere. Intuitively, we could think of the retailer and independent seller agreeing in advance on what prices they will set for each choice of $\alpha$, given common knowledge of their prices $c_{r},c_{s}$. We could also think of the retailer as more powerful and give them the ability to choose $\rho$, requiring that the independent seller agree to choose the prices outlined in $\rho$ as a condition of the revenue sharing program..
To give some concrete examples, some natural choices for $\rho$ are:
  1. (I)
    $\underline{\rho}(\alpha)$ – for each $\alpha$, $\underline{\rho}(\alpha)\in\mathcal{S}_{p}$ is an admissible, Pareto optimal price configuration where the equilibrium outcome has demand being fulfilled at the minimum possible price
  2. (II)
    $\overline{\rho}(\alpha)$ – for each $\alpha$, $\overline{\rho}(\alpha)\in\mathcal{S}_{p}$ is an admissible, Pareto optimal price configuration where the equilibrium outcome has demand being fulfilled at the maximum possible price
By inspection of (26) and (27), the only region in which $\rho$ makes a nontrivial choice is in the gray continuum region of Figure 3, which allows demand to be fulfilled at $p\in\left[\max\left\{p_{r^{*}},p_{s_{\text{ind}}}\right\},\min\left\{p_{s^{*}},p^{\dagger}\right\}\right]$ when $\alpha\in\left[\max\left\{\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\right\},\alpha^{\dagger}_{s}\right]$. Continuing our examples, we would have that
  1. (I)
    For all $\alpha\in\left[\max\left\{\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\right\},\alpha^{\dagger}_{s}\right]$, $\underline{\rho}(\alpha)\equiv\left(\max\left\{p_{r^{*}},p_{s_{\text{ind}}}\right\},\max\left\{p_{r^{*}},p_{s_{\text{ind}}}\right\}\right)\in\mathcal{S}_{p}$
  2. (II)
    For all $\alpha\in\left[\max\left\{\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\right\},\alpha^{\dagger}_{s}\right]$, $\overline{\rho}(\alpha)\equiv\left(\min\left\{p_{s^{*}},p^{\dagger}\right\},\min\left\{p_{s^{*}},p^{\dagger}\right\}\right)\in\mathcal{S}_{p}$
For all $\alpha\notin\left[\max\left\{\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\right\},\alpha^{\dagger}_{s}\right]$, there is only one admissible, Pareto optimal equilibrium outcome as specified by (26) and (27), thus we will always observe the same outcome irrespective of our choice of $\rho$. For any $\rho$, the subgame perfect, admissible, Pareto optimal equilibria are as follows:
Equilibrium of Fee Optimization Game
Given
$\rho(\alpha,c_{r},c_{s})\in\mathcal{S}_{p}$
– an admissible, Pareto optimal equilibrium strategy profile of the shared revenue Bertrand game for every set of parameters
Equilibrium
  1. (1)
    The retailer sets $\alpha_{*}(c_{r},c_{s},\rho)$, the minimum referral fee that maximizes equilibrium payoff
  2. (2)
    The retailer and independent seller set the equilibrium price configuration $\rho(\alpha_{*},c_{r},c_{s})$
5.2. Interpreting equilibria
To complete our analysis, our natural next step is to find $\alpha_{*}$. We’ll denote as $\pi_{r}^{\text{(eq)}}(\rho(\alpha),\alpha,c_{r},c_{s})$ the payoff function in equilibrium of the shared-revenue Bertrand game for the retailer with strategy $\rho$ and $\pi_{s}^{\text{(eq)}}$ the analogous function for the independent seller. With this notation, it’s clear that
(28) $\alpha_{*}(c_{r},c_{s},\rho)\equiv\min_{\alpha}\left[\arg\max_{\alpha}\pi_{r}^{\text{(eq)}}(\rho(\alpha),\alpha,c_{r},c_{s})\right]$
However, the functional dependence of $\alpha_{*}$ on $\rho$ makes this optimization problem complicated in general. To build intuition, we will begin by considering only $\underline{\rho}$ and $\overline{\rho}$. In Appendix F.4, we show the simple result that
$\alpha_{*}(c_{r},c_{s},\underline{\rho})=\alpha_{r^{*}}\equiv\underline{\alpha}$
In other words, if the retailer and independent seller agree to play strategy $\underline{\rho}$, the retailer will always choose referral fee $\underline{\alpha}=\alpha_{r^{*}}$ in equilibrium of the Fee Optimization Game.
The optimal fee for $\overline{\rho}$ is slightly more complicated than for $\underline{\rho}$, but we show in Appendix F.3 that
(29) $\alpha_{*}(c_{r},c_{s},\overline{\rho})=\begin{cases}\arg\max_{\alpha}\pi_{r,s}(p_{s^{*}},\alpha)\equiv\overline{\alpha}&\exists\alpha\text{ s.t. }\pi_{r,s}(p_{s^{*}},\alpha)>\pi_{r,r}(p_{r^{*}})\\ \alpha^{\dagger}_{r}&\text{else}\end{cases}$
Figure 4. Equilibrium payoff curves
To illustrate $\underline{\alpha}$ and $\overline{\alpha}$, we plot typical payoff curves in Figure 4. The background of the plot corresponds to the equilibria in Figure 3. In the left plot, there does exist an $\alpha$ such that $\pi_{r,s}(p_{s^{*}},\alpha)>\pi_{r,r}(p_{r^{*}})$, so $\alpha_{*}(\overline{\rho})=\overline{\alpha}$. In the right plot, for all $\alpha$ we have $\pi_{r,s}(p_{s^{*}},\alpha)\leq\pi_{r,r}(p_{r^{*}})$, so $\alpha_{*}(\overline{\rho})=\alpha^{\dagger}_{r}$.
5.3. General equilibrium strategy profiles
In Section 5.2, we interpreted the equilibria of the Fee Optimization Game under specifically the equilibrium strategy profiles $\underline{\rho}$ and $\overline{\rho}$. However, as a result of monotonicity of the payoff functions in the gray region (the only region where the choice of $\rho$ is nontrivial), we can bound the payoffs for any $\rho$ by the payoffs for $\underline{\rho}$ and $\overline{\rho}$. Thus, the shaded areas in the gray region of Figure 4 correspond to the range of possible payoffs for the retailer and seller, with the exact values depending on the choice of $\rho$. Slightly more formally, in Appendix F.2 we prove the following:
Corollary 0.
For any strategy profile $\rho$, the retailer’s (seller’s) equilibrium payoff is upper bounded (lower bounded) by their equilibrium payoff under $\underline{\rho}$, and the retailer’s (seller’s) equilibrium payoff is lower bounded (upper bounded) by their equilibrium payoff under $\overline{\rho}$.
In light of Corollary 1, we can bound the equilibrium payoff of the Fee Optimization Game $\pi_{r}$ for any strategy profile $\rho$
$\max_{\alpha}\pi_{r}^{\text{(eq)}}(\overline{\rho}(\alpha),\alpha,c_{r},c_{s})\leq\pi_{r}^{\text{(eq)}}(\rho(\alpha_{*}),\alpha_{*},c_{r},c_{s})=\pi_{r}(\rho(\alpha),c_{r},c_{s})\leq\max_{\alpha}\pi_{r}^{\text{(eq)}}(\underline{\rho}(\alpha),\alpha,c_{r},c_{s})$
Bounds on $\pi_{s}$ and $\alpha_{*}$ follow similarly by considering $\pi_{r}^{\text{(eq)}}$ under the strategy profiles $\underline{\rho}$ and $\overline{\rho}$. Indeed, we show in Appendix F.3 and F.4 that for any equilibrium strategy profile $\rho$, the admissible, Pareto optimal, subgame perfect Nash equilibria of the sequential game satisfy:
(30) $\pi_{r}\in\begin{cases}\left[\pi_{r,s}(p_{s^{*}},\overline{\alpha}),\pi_{r,s}\left(p_{r^{*}},\underline{\alpha}\right)\right]&\exists\alpha\text{ s.t. }\pi_{r,s}(p_{s^{*}},\alpha)>\pi_{r,r}(p_{r^{*}})\\ \left[\pi_{r,r}(p_{r^{*}}),\pi_{r,s}\left(p_{r^{*}},\underline{\alpha}\right)\right]&\text{else}\end{cases}$
(31) $\pi_{s}\in\left[0,\pi_{s,s}\left(p_{s^{*}},\min\left\{\alpha_{s^{*}},\alpha^{\dagger}_{r}\right\}\right)\right]\ ;\qquad\alpha_{*}\in\left[\min\left\{\alpha_{s^{*}},\alpha^{\dagger}_{r}\right\},\alpha^{\dagger}_{s}\right]$
One important takeaway from (30) and (31) is that the equilibrium $\alpha_{*}$ is bounded away from $1$. The existence of such an optimal fee is echoed by optimal taxation theories such as the Laffer curve (Laffer, 2004) – it is not in the retailer’s interest to take all the independent seller’s revenue.
In light of this general result, our analysis from Section 5.2 has far-reaching implications. Importantly, examining Figure 4, we see that the retailer will always induce an equilibrium where the independent seller fulfills demand at some $p\in\left[\max\left\{p_{s_{\text{ind}}},p_{r^{*}}\right\},\left\{p_{s^{*}},p^{\dagger}\right\}\right]$ depending on $\rho$ (red or gray regions). Interestingly, the socially suboptimal green region where the independent seller fulfills demand at $p_{r_{\text{ind}}}$ is never observed in equilibrium of the Fee Optimization Game.
We do also find that the retailer never sets the fee high enough such that they fulfill demand themselves at price $p_{r^{*}}$ (blue region). However, this is less interesting, because it is an artifact of the relative Pareto optimality criteria that we applied when refining equilibria in Section 4.2; we argued there that the retailer would prefer for the independent seller to fulfill demand at $p^{\dagger}$ rather fulfilling demand themselves at $p_{r^{*}}$.
6. An outside option for the independent seller
In the Fee Optimization Game, the retailer was free to choose any fee they’d like, and the independent seller would stay in the revenue sharing program irrespective of the fee chosen. However, to make the game more realistic, we should model the independent seller’s choice to enter the revenue sharing program as a negotiation with the retailer. Indeed, in real scenarios the independent seller often has a compelling alternative (outside option) – selling their product independently rather than participating in the revenue sharing program. Formally, we have the following game:
Outside Option Game
Parameters
Retailer’s cost $c_{r}$, seller’s cost $c_{s}$, outside option cost differential $\delta\in(0,c_{r}-c_{s})$
Gameplay
  1. (1)
    The retailer chooses a referral fee $\alpha\in(0,1)$
  2. (2)
    The independent seller chooses one of two options:
    Stay in the retailer’s revenue sharing program
    Leave the retailer’s revenue sharing program and sell on their own
  3. (3)
    The retailer and seller choose their prices:
    If the independent seller chose to stay, they play the shared-revenue Bertrand game
    If the independent seller chose to leave, they play a to-be-defined “leaving” subgame
6.1. Leaving subgame
To finish defining the Outside Option Game, we must outline the structure of this “leaving” subgame. When not clear from context, we will use superscripts $\cdot^{(\ell)}$ to denote relevant quantities in the leaving subgame and $\cdot^{(o)}$ to denote relevant quantities in the outside option game as a whole. There are many reasonable specifications for this leaving subgame, and more sophisticated models may better capture reality, but as a first step we choose a very simple model: if the independent seller chooses to leave, they will play a standard Bertrand game against the retailer.
With the constraint that $\alpha>0$, it seems obvious that the independent seller would always leave to avoid paying the referral fee. Thus, we will posit that if they choose to sell on their own, the effective cost to the independent seller will be $c_{s}+\delta$. This additive $\delta$ factor could represent any number of costs that the independent seller might incur if selling without the retailer such as advertising, shipping, or storing inventory. It could even represent an opportunity cost that the smaller independent seller incurs by not exposing their product to the larger retailer’s customers.
Whatever it represents, we will assume that there exists a $\delta>0$ such that the independent seller does not always prefer to leave. We now have a standard Bertrand game with potentially asymmetric costs depending on the value of $\delta$, for which the equilibria have been derived in (Blume, 2003; Kartik, 2011) for instance. However, if $\delta\geq c_{r}-c_{s}$ such that $c_{s}+\delta\geq c_{r}$, the independent seller achieves an equilibrium payoff of 00 if they exercise their outside option, making leaving the revenue sharing program a non-credible threat. This can be seen by adapting the derivation in Appendix G to one where $c_{r}\leq c_{s}+\delta$ instead of $c_{s}+\delta<c_{r}$. Thus, with the constraint that $\delta\in(0,c_{r}-c_{s})$, we show in Appendix G that the admissible equilibrium outcomes of the leaving subgame are
$(p,x)=\left(\min\left\{c_{r},p_{s^{*}}^{(\ell)}\right\},s\right)$
where $p_{s^{*}}^{(\ell)}\equiv\frac{1}{2}(1+c_{s}+\delta)$ is the maximizer of the independent seller’s payoff function in the leaving subgame. Importantly, because the retailer never fulfills demand in equilibrium and there is no revenue sharing, they always achieve an equilibrium payoff of 00 if the independent seller chooses to leave. Because the retailer achieves nonzero payoff in equilibrium otherwise, they would always prefer to choose an $\alpha$ such that the independent seller stays over an $\alpha$ such that the independent seller leaves. Furthermore, notice that the leaving subgame itself is entirely independent of $\alpha$.
6.2. Refined equilibria of outside option game
The analysis in Appendix F.1 about refinements for the Fee Optimization Game equilibria applies to the Outside Option Game equilibria: a strategy of the sequential game is admissible if $\rho$ is admissible, and Pareto optimality implies that the retailer will choose the lowest fee possible to achieve their maximum payoff. Furthermore, because the retailer achieves nonzero payoff if the independent seller chooses to stay and 00 payoff if they leave, Pareto optimality also tells us that if the independent seller is indifferent between staying and leaving, they will choose to stay. To summarize, the admissible, Pareto optimal, subgame perfect equilibrium strategies are:
Equilibrium of Outside Option Game
Given
$\rho(\alpha,c_{r},c_{s})\in\mathcal{S}_{p}$
– an admissible, Pareto optimal equilibrium strategy profile of the shared-revenue Bertrand game for every set of parameters
Equilibrium
  1. (1)
    The retailer sets $\alpha^{(o)}_{*}(\delta,c_{r},c_{s},\rho)$, the minimum referral fee that maximizes equilibrium payoff and does not cause the independent seller to strictly prefer leaving
  2. (2)
    The independent seller stays in the revenue sharing program
  3. (3)
    The retailer and independent seller play the shared-revenue Bertrand game and simultaneously set the equilibrium price $\rho(\alpha^{(o)}_{*},c_{r},c_{s})$
To be precise, $\alpha^{(o)}_{*}$ must satisfy the following property similar to (28):
(32) $\alpha^{(o)}_{*}(\delta,c_{r},c_{s},\rho)\equiv\min_{\alpha}\left[\arg\max_{\alpha}\left\{\pi_{r}^{\text{(eq)}}(\alpha,\rho)\mid\pi_{s}^{\text{(eq)}}(\alpha,\rho)\geq\pi_{s}^{(\ell)}(\delta)\right\}\right]$
Figure 5. Equilibrium payoff curves with an outside option
As before, because of the functional dependence of $\alpha^{(o)}_{*}$ on $\rho$, this is complicated in general. Just as in Section 5.2, we’ll build intuition by analyzing simpler cases first. In Figure 5, we plot the same equilibrium payoff curves as Figure 4, however we now include an outside option with $\delta=0.062$ in the left subplot and $\delta=0.003$ in the right subplot. For simplicity, in the left subplot we assume that $\rho=\overline{\rho}$. If the retailer sets the fee above some threshold, the independent seller would prefer to leave the revenue sharing program and sell on their own instead. It will be helpful to give this threshold a name; let’s define
(33) $\alpha_{\max}(\delta,c_{r},c_{s},\rho)\equiv\sup\left\{\alpha\mid\pi_{s}^{\text{(eq)}}(\alpha,\rho)\geq\pi_{s}^{(\ell)}(\delta)\right\}$
If the retailer chooses a fee $\alpha>\alpha_{\max}$, they will achieve a payoff of 00 in equilibrium of the Outside Option Game because the independent seller will leave. The independent seller, on the other hand, will achieve a payoff of $\pi_{s}^{(\ell)}(\delta)$ for $\alpha>\alpha_{\max}$, which is therefore a lower bound on their equilibrium payoff. Additionally, in equilibrium of the Outside Option Game the retailer will choose a fee $\alpha\leq\alpha_{\max}$, which bounds the maximum fee further away from $1$.
Interestingly, the left subplot of Figure 5 demonstrates that the retailer will not necessarily choose $\alpha^{(o)}_{*}=\alpha_{\max}$, indeed their payoff under $\overline{\rho}$ is maximized by choosing $\overline{\alpha}\in(0,\alpha_{\max})$. However, if we simply switch the equilibrium strategy profile from $\overline{\rho}$ to $\underline{\rho}$, the retailer will choose $\alpha^{(o)}_{*}=\alpha_{\max}$, which serves as a good reminder that the choice of $\rho$ may have a significant effect on the equilibrium outcome in general.
The right subplot of Figure 5 shows another interesting case where $\delta$ is low enough such that the independent seller achieves a higher payoff than the retailer in equilibrium. In fact, the independent seller is so competitive with the retailer that the retailer achieves a significantly lower payoff in equilibrium of the Outside Option Game than they would if the indepdendent seller did not exist. Additionally, notice that $\alpha_{\max}<\alpha^{\dagger}_{r}$ is not in the continuum region. Because of this, the choice of $\rho$ does not matter, which should be clear graphically because the entire continuum region is “grayed out” for both the retailer and independent seller. In fact, we can say more generally that
(34) $\alpha^{(o)}_{*}=\min\left\{\alpha_{*},\alpha_{\max}\right\}\quad\text{ if }\alpha_{\max}\leq\max\left\{\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\right\}$
which is a simple special case of (32).
The Outside Option Game is the final extension of the shared-revenue Bertrand game that we explore in this paper. The game captures the effect that the cost configurations ($c_{r},c_{s}$) and the credibility of the independent seller’s outside option threat ($\delta$) have on the optimal fee structures and the minimum/maximum payoffs achievable in equilibrium of the revenue sharing program.
7. Conclusion and future work
In this work, we presented a game theory model of a shared-revenue Bertrand game to analyze the dynamics and incentives induced by revenue sharing programs. Furthermore, we investigated some natural extensions of the model to make it more realistic by modeling the proportion of revenue sharing as an action and giving the independent seller an outside option to sell on their own. Our findings resonate strongly with existing results from duopolistic competition and optimal taxation theory. In equilibrium, both players are able to achieve positive payoff, and the referral fee is set at a “sweet spot” low enough to stimulate economic activity while high enough to maximize payoff.
However, this work is just a starting point, and there are still many future research directions in which this model can be extended with even more potential implications for understanding the dynamics of revenue sharing:
  • To improve the present analysis of the shared-revenue Bertrand game, some natural next steps would be to more completely consider general demand curves and mixed strategies. While we anticipate qualitatively similar equilibria under appropriate assumptions on the demand curves and the mixing distributions, there is much left to formally prove this.
  • We have assumed throughout that this is a game of perfect, complete information – modeling uncertainty or forecasting of the costs or demand curves are interesting next steps.
  • We do not closely consider the manufacturing of the product, but we would expect the dynamics to differ in the cases that (a) there is one primary manufacturer of the product, from whom both the retailer and indepedent seller buy, or (b) there are many manufacturers of the product, in which case the margins are low.
  • We could impose capacity constraints on the players, allowing us to reason about strategies for stocking both retailers’ and independent sellers’ products in the same warehouse.
  • We look only at a duopolistic interaction, but more sophisticated models of entry and exit from a shared-revenue Bertrand marketplace may help make the game more realistic.
  • A large retailer or ridesharing service may not want to choose a different fee for each individual independent seller on its platform, instead preferring to choose an “aggregate fee” for a homogeneous group of independent sellers.
  • Empirical studies of revenue sharing programs could reveal aspects of the real world our models fail to capture, including in the other application domains introduced in Section 1.1.
Ultimately, this work marks an important first step towards understanding cooperative business practices between economic agents from first principles.
Acknowledgements.
We would like to thank Dirk Bergemann for his insightful comments and suggestions.
Appendix A Notation table
Table 1. Notation table
Notation Value(s) Description
$c_{x}$ $\in(0,1)$ Marginal cost for player $x$
$p_{x}$ $\in(0,1)$ Price set by player $x$
$\alpha$ $\in(0,1)$ Referral fee
$\beta$ $\in[0,1]$ Market split proportion if prices are equal
$q_{x}(p_{r},p_{s})$ (5) Quantity demanded for player $x$
$\mathcal{S}_{p}$ $(0,1)\times(0,1)$ Space of price configurations, eg. $(p_{r},p_{s})\in\mathcal{S}_{p}$
$(p,x)$ $\in(0,1)\times\{r,s\}$ Outcome where player $x$ fulfills demand at price $p$
$\delta$ $\in(0,c_{r}-c_{s})$ Additional marginal cost if independent seller leaves
$\rho(\alpha,c_{r},c_{s})$ $\in\mathcal{S}_{p}$ Equilibrium strategy profile of shared-revenue Bertrand game
$\overline{\rho}(\alpha,c_{r},c_{s})$ $\in\mathcal{S}_{p}$ Equilibrium strategy profile with maximum prices
$\underline{\rho}(\alpha,c_{r},c_{s})$ $\in\mathcal{S}_{p}$ Equilibrium strategy profile with minimium prices
$\pi_{x}(\cdot)$ $\in[0,1)$ Player $x$’s payoff
$\pi_{x}^{\text{(eq)}}(\rho,\alpha,c_{r},c_{s})$ $\in[0,1)$ Payoff in equilibrium of shared-revenue Bertrand game
$\pi_{x}^{(\ell)}(\delta,c_{r},c_{s})$ $\in[0,1)$ Payoff in equilibrium of leaving subgame
$\pi_{x,y}(p_{y},\alpha)$ $\in[0,1)$ Player $x$’s payoff when $y$ fulfills demand at $p_{y}$
$c_{s^{*}}$ $\frac{2c_{r}^{2}}{1+c_{r}}$ Seller optimal feasibility cost; $c_{s}\leq c_{s^{*}}$ is necessary for $(p_{s^{*}},s)$ to be a potential equilibrium
$p_{s^{*}}$ $\frac{1}{2}\left(1+\frac{c_{s}}{1-\alpha}\right)$ Seller optimal single agent price, the maximizer of $\pi_{s,s}$
$p_{s^{*}}^{(\ell)}$ $\frac{1}{2}\left(1+c_{s}+\delta\right)$ Seller optimal single agent price in leaving subgame
$p_{r^{*}}$ $\frac{1}{2}\left(1+c_{r}\right)$ Retailer optimal single agent price, the maximizer of $\pi_{r,r}$
$p_{r_{\text{ind}}}$ $\frac{c_{r}}{1-\alpha}$ Retailer indifference price, satisfying $\pi_{r,r}(p_{r_{\text{ind}}})=\pi_{r,s}(p_{r_{\text{ind}}})$
$p_{s_{\text{ind}}}$ $\frac{c_{s}}{1-\alpha}$ Seller indifference price, which is also their breakeven price as it satisfies $\pi_{s,s}(p_{s_{\text{ind}}})=\pi_{s,r}(p_{s_{\text{ind}}})=0$
$p^{\dagger}$ (12) Retailer optimality equivalence price, satisfying $\pi_{r,s}(p^{\dagger})=\pi_{r,r}(p_{r^{*}})$
$\alpha^{\dagger}_{s}$ (62) Seller $p^{\dagger}$ relevance fee, below which selling at $p^{\dagger}$ becomes a potential best response
$\alpha^{\dagger}_{r}$ $\frac{1-c_{r}}{1+c_{r}}$ Retailer $p^{\dagger}$ relevance fee, above which allowing the independent seller to sell at $p^{\dagger}$ becomes a potential best response
$\alpha_{s^{*}}$ $1-2c_{r}+c_{s}$ Seller optimal price feasibility fee, above which $(p_{s^{*}},s)$ becomes a potential equilibrium
$\alpha_{r^{*}}$ $1-\frac{2c_{s}}{1+c_{r}}$ Retailer optimal price feasibility fee, above which $(p_{r^{*}},r)$ becomes a potential equilibrium
$\alpha_{s^{*},r^{*}}$ $\frac{c_{r}-c_{s}}{c_{r}}$ Optimality switching fee satisfying $p_{s^{*}}(\alpha_{s^{*},r^{*}})=p_{r^{*}}$
$\alpha_{*}$ (28) Retailer equilibrium alpha in the Fee Optimization Game
$\overline{\alpha}$ See Lemma 4 Retailer equilibrium fee with strategy profile $\overline{\rho}$
$\underline{\alpha}$ $\alpha_{r^{*}}$ Retailer equilibrium fee with strategy profile $\underline{\rho}$
$\alpha^{(o)}_{*}$ (32) Retailer equilibrium fee for the Outside Option Game
$\alpha_{\max}$ (33) Maximum fee for which the independent seller prefers to stay in the revenue sharing program rather than exercise their outside option in the Outside Option Game
Table LABEL:tab:notation summarizes the symbols used throughout the paper along with their value or range of possible values. We use $r$ to denote the retailer and $s$ to denote the seller, with $x,y\in\{r,s\}$ placeholders in quantities that are defined analogously for both players. When clear from context throughout the paper, we drop some of the explicit arguments of the functions – note for instance that the prices and fees implicitly depend on the costs.
Appendix B Justification of game setup
B.1. The retailer and seller can achieve nonzero payoff
We assume in Section 2 that $c_{s}<c_{r}<1$. Indeed, suppose on the other hand that $c_{r}\geq 1$. Then, the game becomes trivial:
Lemma 0.
In a shared-revenue Bertrand game with $c_{r}\geq 1$, the only admissible Nash equilibrium is $(p_{r}=1,p_{s}=\min\left\{p_{s^{*}},1\right\})$.
Proof.
The retailer’s dominant strategy is setting $p_{r}=1$. Consider a price $p_{r}<1$. If $p_{s}\leq p_{r}$,
$\pi_{r}(p_{r},p_{s})=\pi_{r,s}(p_{s})=\pi_{r}(1,p_{s})$
On the other hand, if $p_{s}\in(p_{r},1]$, we have
$\pi_{r}(p_{r},p_{s})=\pi_{r,r}(p_{r})<0<\pi_{r,s}(p_{s})=\pi_{r}(1,p_{s})$
With this, the independent seller’s best response is clearly $p_{s}=\min\left\{1,p_{s^{*}}\right\}$. ∎
Enforcing this requirement allows us to avoid such trivial equilibria where one or both players abstain from the market because they will simply never fulfill demand at any price.
B.2. Independent seller’s cost is lower
We also assume in Section 2 that $c_{s}<c_{r}$. Suppose on the other hand that $c_{r}\leq c_{s}$. Then, the game becomes trivial:
Lemma 0.
In a shared-revenue Bertrand game with $c_{r}<c_{s}$, there are no Nash equilibria where the independent seller fulfills demand.
Proof.
Suppose for contradiction there exists an equilibrium $(p_{r}\geq p_{s},p_{s})$ where the independent seller fulfills demand. Then, we claim the following must be satisfied:
  1. (I)
    $p_{s}\geq p_{s_{\text{ind}}}$
  2. (II)
    $p_{s}\leq p_{r_{\text{ind}}}$
Indeed, to show (I) suppose for contradiction $p_{s}\leq p_{r}<p_{s_{\text{ind}}}$. Then, we have
$\pi_{s}(p_{r},p_{s})=\pi_{s,s}(p_{s})<0=\pi_{s,r}(p_{r})=\pi_{s}(p_{r},p_{s_{\text{ind}}})$
Similarly, if $p_{s}<p_{s_{\text{ind}}}\leq p_{r}$, we have $\pi_{s}(p_{r},p_{s_{\text{ind}}})=\pi_{s,s}(p_{s_{\text{ind}}})=0$.
To show (II), suppose for contradiction $p_{r_{\text{ind}}}<p_{s}\leq p_{r}$. Then, by construction of $p_{r_{\text{ind}}}$, we have
$\pi_{r}(p_{r},p_{s})=\pi_{r,s}(p_{s})<\pi_{r,r}(p_{s})$
by continuity of $\pi_{r,r}$, this implies that there exists some $\varepsilon>0$ such that $\pi_{r,r}(p_{s}-\varepsilon)>\pi_{r,s}(p_{s})$, and therefore $p_{r}$ yields a lower payoff than $p_{s}-\varepsilon$ so the retailer is not best responding.
However, both (I) and (II) cannot be simultaneously satisfied, because under the assumption that $c_{r}<c_{s}$ we have that
$p_{r_{\text{ind}}}=\frac{c_{r}}{1-\alpha}<\frac{c_{s}}{1-\alpha}=p_{s_{\text{ind}}}$
Thus, such an equilibrium cannot exist. ∎
Because we are interested in modeling the dynamics between the retailer and independent seller, if there are no equilibria where the independent seller fulfills demand our model is entirely uninteresting. Intuitively, if the retailer can fulfill demand themselves for less cost per unit, they have no incentive to initiate a revenue sharing program in the first place.
This implies that the revenue sharing program is best thought of as permitting the sale of niche items or services that the retailer does not specialize in. Instead of specializing in this offering themselves, the model aims to find the conditions under which it would be beneficial for the retailer to partner with the specialist independent seller.
Appendix C Proof of Proposition 1
Proposition 0.
Let $(p_{r},p_{s})=(p,p)$, with the equilibrium quantities sold by each player $q_{r}=\beta(1-p)$ and $q_{s}=(1-\beta)(1-p)$. Then $\beta\in\{0,1\}$.
Proof.
Let $p$ be an equilibrium market price, with the equilibrium quantities sold by each player $q_{r}=\beta(1-p)$ and $q_{s}=(1-\beta)(1-p)$, with $\beta\notin\{0,1\}$. Then the total quantity sold is $q=q_{r}+q_{s}=1-p$ and the retailer’s payoff is
(35) $\pi_{r}$ $=(p-c_{r})q_{r}+\alpha pq_{s}$
$=(p-c_{r})\beta(1-p)+\alpha p(1-\beta)(1-p)$
$=\left[(p-c_{r})\beta+\alpha p(1-\beta)\right](1-p).$
We examine three possible cases — $(p-c_{r})>\alpha p$, $(p-c_{r})<\alpha p$ and $(p-c_{r})=\alpha p$ — and show that in each case, one of the players has an incentive to deviate to a different price.
If $(p-c_{r})>\alpha p$ (the retailer’s payoff per unit it sells is larger than its payoff from the independent seller fullfilling that unit), then the retailer prefers to deviate by setting a price $p-\varepsilon$ to take the entire market. Similarly, if $(p-c_{r})<\alpha p$, the retailer can achieve a higher payoff by deviating to price $p+\varepsilon$, allowing the the independent seller to fulfill the demand. When $(p-c_{r})=\alpha p$, the independent seller’s payoff is $q_{s}=(1-\beta)(p-c_{s}-\alpha p)$. A similar argument, comparing $(p-c_{s}-\alpha p)$ and 00, shows that the independent seller will either decrease its price to take the entire market, or increase its price and allow the retailer to fulfill the demand (since the independent seller only achieves nonzero payoff when they fulfills demand themselves, this last case will only happen if $p$ is below the independent seller’s breakeven price). Note that the case of $(p-c_{r})=\alpha p$ and $(p-c_{s}-\alpha p)=0$ is impossible, as that would imply $c_{r}=c_{s}$, which contradicts our assumption that $c_{r}>c_{s}$. Thus, if $\beta\notin\{0,1\}$, the proposed price and market split does not constitute an equilibrium. ∎
Appendix D Detailed equilibria derivation
This section is dedicated to deriving the Nash equilibria of (22) and (23). For clarity of exposition, we outline conditions that any Nash equilibrium price configuration $(p_{r},p_{s})\in\mathcal{S}_{p}$ must satisfy. By definition of a Nash equilibrium, the following must hold:
  1. (I)
    $\sup_{p_{r}^{\prime}<p_{s}}\pi_{r}(p_{r}^{\prime},p_{s})\leq\pi_{r}(p_{r},p_{s})$
  2. (II)
    $\sup_{p_{r}^{\prime}>p_{s}}\pi_{r}(p_{r}^{\prime},p_{s})\leq\pi_{r}(p_{r},p_{s})$
  3. (III)
    $\sup_{p_{s}^{\prime}<p_{r}}\pi_{s}(p_{r},p_{s}^{\prime})\leq\pi_{s}(p_{r},p_{s})$
  4. (IV)
    $\sup_{p_{s}^{\prime}>p_{r}}\pi_{s}(p_{r},p_{s}^{\prime})\leq\pi_{s}(p_{r},p_{s})$
D.1. Shared-price equilibria $p_{r}=p=p_{s}$ in which the retailer fulfills demand
Condition IV always holds, as the independent seller achieves a payoff of zero whether they set some price $p^{\prime}>p$ or they set price $p$ and allow the retailer to fulfill all demand.
Condition I requires the retailer would not prefer to set a price $p^{\prime}<p$. Because $\pi_{r,r}$ is a concave quadratic function, it is increasing on $[0,p_{r^{*}}]$, so condition I holds if and only if $p\leq p_{r^{*}}$. Indeed, we derive an expression for $p_{r^{*}}$ by differentiating the quadratic $\pi_{r,r}$ and setting it equal to 00:
$\frac{\partial\pi_{r,r}}{\partial p_{r}}$ $=\frac{\partial}{\partial p_{r}}(p_{r}-c_{r})(1-p_{r})$
$=-2p_{r}+1+c_{r}\stackrel{{\scriptstyle\text{set}}}{{=}}0$
(36) $\implies p_{r^{*}}$ $=\frac{1+c_{r}}{2}$
Condition II is equivalent to saying that the retailer would not prefer to increase their price, effectively allowing the independent seller to fulfill demand at price $p$ instead of the retailer fulfilling demand at price $p$. Formally, we need
(37) $\pi_{r,s}(p)$ $\leq\pi_{r,r}(p)$
$\alpha p(1-p)$ $\leq(p-c_{r})(1-p)$
$\alpha p$ $\leq(p-c_{r})$
$p_{r_{\text{ind}}}=\frac{c_{r}}{1-\alpha}$ $\leq p$
Indeed, $p_{r_{\text{ind}}}$ is defined as the price at which the retailer’s payoff is equal no matter which seller fulfills demand, which is exactly equal to $\frac{c_{r}}{1-\alpha}$
$\pi_{r,r}(p_{r_{\text{ind}}})$ $=\pi_{r,s}(p_{r_{\text{ind}}})$
$(p_{r_{\text{ind}}}-c_{r})(1-p_{r_{\text{ind}}})$ $=\alpha p_{r_{\text{ind}}}(1-p_{r_{\text{ind}}})$
$-c_{r}$ $=\alpha p_{r_{\text{ind}}}-p_{r_{\text{ind}}}$
(38) $p_{r_{\text{ind}}}$ $=\frac{c_{r}}{1-\alpha}$
With this intuition, the simplification in (37) is almost true by definition. All that is required is to show the direction of the inequality, which is correct because the retailer’s payoff from the independent seller fulfilling demand at $p=0$ is zero, but is negative when it fulfills demand on its own at $p=0$.
Condition III posits that the independent seller would not prefer to undercut the retailer and fulfill demand themselves at some price $p^{\prime}<p$. However, recall that in an equilibrium where the retailer fulfills demand, the independent seller achieves a payoff of 00. Thus, we can rewrite condition III as:
(39) $\sup_{p^{\prime}\in[0,p)}\left[\pi_{s,s}(p^{\prime})\right]$ $\leq\pi_{s,r}(p)$
$\sup_{p^{\prime}\in[0,p)}\left[\pi_{s,s}(p^{\prime})\right]$ $\leq 0$
Because $\pi_{s,s}$ is a concave quadratic function, we have that
$\sup_{p^{\prime}\in[0,p)}\left[\pi_{s,s}(p^{\prime})\right]=\begin{cases}\pi_{s,s}(p)&p\leq p_{s^{*}}\\ \pi_{s,s}(p_{s^{*}})&p>p_{s^{*}}\end{cases}$
Where $p_{s^{*}}$ is the maximum of $\pi_{s,s}$, for which we find a closed-form –
$\frac{\partial\pi_{s,s}}{\partial p_{s}}$ $=\frac{\partial}{\partial p_{s}}((1-\alpha)p_{s}-c_{s})(1-p_{s})$
$=-2(1-\alpha)p_{s}+(1-\alpha)+c_{s}\stackrel{{\scriptstyle\text{set}}}{{=}}0$
(40) $\implies p_{s^{*}}$ $=\frac{1+\frac{c_{s}}{1-\alpha}}{2}$
If $p_{s^{*}}<p<1$, then $0<\pi_{s,s}(p)\leq\sup_{p^{\prime}\in[0,p)}\left[\pi_{s,s}(p^{\prime})\right]$ and the condition never holds. On the other hand, if $p\leq p_{s^{*}}$, then (39) reduces to
(41) $\pi_{s,s}(p)$ $\leq 0$
$((1-\alpha)p-c_{s})(1-p)$ $\leq 0$
$p$ $\leq\frac{c_{s}}{1-\alpha}=p_{s_{\text{ind}}}$
where we ignored the $1\leq p$ because $p\in[0,1]$. Indeed, we have that $p_{s_{\text{ind}}}=\frac{c_{s}}{1-\alpha}$ because
$\pi_{s,s}(p_{s_{\text{ind}}})$ $=\pi_{s,r}(p_{s_{\text{ind}}})$
$((1-\alpha)p_{s_{\text{ind}}}-c_{s})(1-p_{s_{\text{ind}}})$ $=0$
(42) $p_{s_{\text{ind}}}$ $=\frac{c_{s}}{1-\alpha}$
Combining the conditions, we have shown that an equilibrium price in which the retailer fulfills demand must satisfy $p\in\left[\max\{0,p_{r_{\text{ind}}}\},\min\{1,p_{s_{\text{ind}}},p_{r^{*}},p_{s^{*}}\}\right]$. However, because $c_{s}<c_{r}$, we have
(43) $p_{s_{\text{ind}}}=\frac{c_{s}}{1-\alpha}<\frac{c_{r}}{1-\alpha}=p_{r_{\text{ind}}}$
so the interval is always empty and there exist no equilibria in this case.
D.2. Shared-price equilibria $p_{r}=p=p_{s}$ in which the independent seller fulfills demand
Condition II always holds, as the retailer does not fulfill any demand whether they set some price $p^{\prime}>p$ or price $p$ and thus achieve the same payoff either way.
Condition III requires that the independent seller would not prefer to fulfill demand at $p^{\prime}<p$. Because $\pi_{s,s}$ is a concave quadratic function maximized at $p_{s^{*}}$, condition III is equivalent to $p\leq p_{s^{*}}$.
Condition IV requires that the independent seller would not prefer that the retailer fulfill demand; it simplifies to
(44) $\pi_{s,r}(p)$ $\leq\pi_{s,s}(p)$
00 0 $\leq((1-\alpha)p-c_{s})(1-p)$
$p_{s_{\text{ind}}}=\frac{c_{s}}{1-\alpha}$ $\leq p$
Finally, condition I posits that the retailer would not prefer to undercut the independent seller and fulfill demand themselves at some price $p^{\prime}<p$. Formally, we can rewrite condition I as
$\sup_{p^{\prime}\in[0,p)}\left[\pi_{r,r}(p^{\prime})\right]\leq\pi_{r,s}(p)$
Because $\pi_{r,r}$ is a concave quadratic function, we have that
(45) $\sup_{p^{\prime}\in[0,p)}\left[\pi_{r,r}(p^{\prime})\right]=\begin{cases}\pi_{r,r}(p)&p\leq p_{r^{*}}\\ \pi_{r,r}(p_{r^{*}})&p>p_{r^{*}}\end{cases}$
D.2.1. Equilibria where $p\leq p_{r^{*}}$
Consider the first case of (45) where $\sup_{p^{\prime}\in[0,p)}\left[\pi_{r,r}(p^{\prime})\right]=\pi_{r,r}(p)$. Condition I then simply becomes $\pi_{r,r}(p)\leq\pi_{r,s}(p)$, which we showed in (37) is equivalent to $p\leq p_{r_{\text{ind}}}$. Thus, we have found a continuum of equilibria $p\in\left[\max\{0,p_{s_{\text{ind}}}\},\min\{1,p_{r_{\text{ind}}},p_{r^{*}},p_{s^{*}}\}\right]$, so long as the interval is nonempty. Note that $p_{s_{\text{ind}}}\geq 0$ and $p_{r^{*}}\leq 1$ always, so we can drop those.
Conditions under which the interval is nonempty.
We showed in (43) that $p_{s_{\text{ind}}}<p_{r_{\text{ind}}}$ always, and we have the following condition for $p_{s_{\text{ind}}}\leq p_{r^{*}}$
$p_{s_{\text{ind}}}$ $\leq p_{r^{*}}$
$\frac{c_{s}}{1-\alpha}$ $\leq\frac{1+c_{r}}{2}$
$2c_{s}$ $\leq(1-\alpha)(1+c_{r})$
(46) $\alpha$ $\leq 1-\frac{2c_{s}}{1+c_{r}}=\alpha_{r^{*}}$
Where $\alpha_{r^{*}}$ gets its name because $p_{r^{*}}\geq p_{s_{\text{ind}}}\iff\alpha\geq\alpha_{r^{*}}$ is necessary for $(p_{r^{*}},r)$ to be an equilibrium. Furthermore, we can see that $p_{s_{\text{ind}}}\leq p_{s^{*}}$ is equivalent to the following –
$p_{s_{\text{ind}}}$ $\leq p_{s^{*}}$
$\frac{c_{s}}{1-\alpha}$ $\leq\frac{1+\frac{c_{s}}{1-\alpha}}{2}$
$2c_{s}$ $\leq 1-\alpha+c_{s}$
(47) $\alpha$ $\leq 1-c_{s}$
However, this is implied by $\alpha\leq\alpha_{r^{*}}$; we can see easily that $\alpha_{r^{*}}\leq 1-c_{s}$ because
(48) $\alpha_{r^{*}}=1-\frac{2c_{s}}{1+c_{r}}$ $\leq 1-c_{s}$
$c_{s}$ $\leq\frac{2c_{s}}{1+c_{r}}$
$c_{r}$ $\leq 1$
This intuitively makes sense, because as we argued in (9), the independent seller cannot attain positive payoff if $\alpha\geq 1-c_{s}$, so of course in this limit the equilibrium $(p_{r^{*}},r)$ is feasible. Thus the interval of prices is nonempty if and only if $\alpha\leq\alpha_{r^{*}}$.
Conditions under which $p_{r_{\text{ind}}}$ is the upper bound.
We have the following condition for $p_{r_{\text{ind}}}\leq p_{r^{*}}$
$p_{r_{\text{ind}}}$ $\leq p_{r^{*}}$
$\frac{c_{r}}{1-\alpha}$ $\leq\frac{1+c_{r}}{2}$
$2c_{r}$ $\leq 1+c_{r}-\alpha-\alpha c_{r}$
$\alpha(1+c_{r})$ $\leq 1-c_{r}$
(49) $\alpha$ $\leq\frac{1-c_{r}}{1+c_{r}}=\alpha^{\dagger}_{r}$
It will become clearer in Appendix D.2.2 where the $\alpha^{\dagger}_{r}$ notation comes from. Furthermore, $p_{r_{\text{ind}}}\leq p_{s^{*}}$ when
$p_{r_{\text{ind}}}$ $\leq p_{s^{*}}$
$\frac{c_{r}}{1-\alpha}$ $\leq\frac{1+\frac{c_{s}}{1-\alpha}}{2}$
$2c_{r}$ $\leq 1-\alpha+c_{s}$
(50) $\alpha$ $\leq 1-2c_{r}+c_{s}=\alpha_{s^{*}}$
Note that $\alpha^{\dagger}_{r}\leq\alpha_{r^{*}}$ always because
(51) $\alpha^{\dagger}_{r}=\frac{1-c_{r}}{1+c_{r}}$ $<1-\frac{2c_{s}}{1+c_{r}}=\alpha_{r^{*}}$
$1-c_{r}$ $<1+c_{r}-2c_{s}$
$c_{s}$ $<c_{r}$
We acknowledge that it is not yet clear when $\alpha_{s^{*}}\stackrel{{\scriptstyle?}}{{\leq}}\alpha^{\dagger}_{r}$ or $\alpha_{s^{*}}\stackrel{{\scriptstyle?}}{{\leq}}\alpha_{r^{*}}$; this is a question to which we will return shortly. However, we can conclude for now that $p_{r_{\text{ind}}}$ is the upper bound when $\alpha\leq\min\left\{\alpha_{s^{*}},\alpha^{\dagger}_{r}\right\}$.
Conditions under which $p_{s^{*}}$ is the upper bound.
We see that $p_{s^{*}}\leq p_{r^{*}}$ when
$p_{s^{*}}$ $\leq p_{r^{*}}$
$(1-\alpha)c_{r}$ $\geq c_{s}$
(52) $\alpha$ $\leq 1-\frac{c_{s}}{c_{r}}=\frac{c_{r}-c_{s}}{c_{r}}=\alpha_{s^{*},r^{*}}$
The notation $\alpha_{s^{*},r^{*}}$ comes from the implication of (52) between optimal prices for retailer and seller. Additionally, per (50), we have that $p_{s^{*}}\leq p_{r_{\text{ind}}}\iff\alpha\geq\alpha_{s^{*}}$, which is exactly the namesake of $\alpha_{s^{*}}$. Furthermore, we can see that $\alpha_{s^{*},r^{*}}\leq\alpha_{r^{*}}$
(53) $\alpha_{s^{*},r^{*}}=\frac{c_{r}-c_{s}}{c_{r}}=1-\frac{c_{s}}{c_{r}}$ $\leq 1-\frac{2c_{s}}{1+c_{r}}=\alpha_{r^{*}}$
This should once again be intuitive because clearly, $(p_{r^{*}},r)$ could not be an equilibrium if $p_{s^{*}}\leq p_{r^{*}}$, as the independent seller would prefer to undercut the retailer and sell at $p_{s^{*}}$ instead. Thus, we have that $p_{s^{*}}$ is the upper bound of the interval when $\alpha\in\left[\alpha_{s^{*}},\alpha_{s^{*},r^{*}}\right]$, provided the interval is nonempty.
Conditions under which $p_{r^{*}}$ is the upper bound.
Combining (52) and (49), we can immediately conclude that $p_{r^{*}}$ is the upper bound of the interval whenever $\alpha\in\left[\max\left\{\alpha_{s^{*},r^{*}},\alpha^{\dagger}_{r}\right\},\alpha_{r^{*}}\right]$.
Conclusion and summary of equilibria.
Explicitly simplifying the condition $\alpha_{s^{*}}\leq\alpha^{\dagger}_{r}$ is uninsightful, so instead we will break up the equilibria into two cases based on which direction of this inequality is true. However, we will prove one equivalence that will make our presentation easier throughout –
$\alpha_{s^{*}}=1-2c_{r}+c_{s}$ $\leq\frac{1-c_{r}}{1+c_{r}}=\alpha^{\dagger}_{r}$
$1-c_{r}-(c_{r}-c_{s})$ $\leq\frac{1-c_{r}}{1+c_{r}}$
$1-\frac{c_{r}-c_{s}}{1-c_{r}}$ $\leq\frac{1}{1+c_{r}}$
$\frac{c_{r}}{1+c_{r}}$ $\leq\frac{c_{r}-c_{s}}{1-c_{r}}$
(54) $\alpha^{\dagger}_{r}=\frac{1-c_{r}}{1+c_{r}}$ $\leq\frac{c_{r}-c_{s}}{c_{r}}=\alpha_{s^{*},r^{*}}$
Note we showed in (51) that $\alpha_{s^{*},r^{*}}\leq\alpha_{r^{*}}$, so as a corollary of (54), we know that $\alpha_{s^{*}}\leq\alpha^{\dagger}_{r}$ implies that $\alpha_{s^{*}}\leq\alpha_{r^{*}}$ as we wanted before. Finally, for ease of presentation, we will manipulate (54) to find an equivalent condition in terms of $c_{s}$ and $c_{s^{*}}$
$\alpha^{\dagger}_{r}=\frac{1-c_{r}}{1+c_{r}}$ $\leq 1-\frac{c_{s}}{c_{r}}=\frac{c_{r}-c_{s}}{c_{r}}=\alpha_{s^{*},r^{*}}$
$\frac{c_{s}}{c_{r}}$ $\leq\frac{2c_{r}}{1+c_{r}}$
(55) $c_{s}$ $\leq\frac{2c_{r}^{2}}{1+c_{r}}=c_{s^{*}}$
To summarize, we have so far found the following equilibria:
Case (i), $c_{s}\leq c_{s^{*}}$:
(56) $p\in\begin{cases}\left[p_{s_{\text{ind}}},p_{r_{\text{ind}}}\right]&\alpha\leq\alpha_{s^{*}}\\ \left[p_{s_{\text{ind}}},p_{s^{*}}\right]&\alpha_{s^{*}}\leq\alpha\leq\alpha_{s^{*},r^{*}}\\ \left[p_{s_{\text{ind}}},p_{r^{*}}\right]&\alpha_{s^{*},r^{*}}\leq\alpha\leq\alpha_{r^{*}}\\ \end{cases}$
Case (ii), $c_{s}\geq c_{s^{*}}$:
(57) $p\in\begin{cases}\left[p_{s_{\text{ind}}},p_{r_{\text{ind}}}\right]&\alpha\leq\alpha^{\dagger}_{r}\\ \left[p_{s_{\text{ind}}},p_{r^{*}}\right]&\alpha^{\dagger}_{r}\leq\alpha\leq\alpha_{r^{*}}\\ \end{cases}$
As we can see from (56) and (57), there only exists an equilibrium $(p_{s^{*}},s)$ when $c_{s}\leq c_{s^{*}}$, which is indeed where the $c_{s^{*}}$ notation comes from.
D.2.2. Equilibria where $p\geq p_{r^{*}}$
Now, consider the second case of (45) where $\sup_{p^{\prime}\in[0,p)}\left[\pi_{r,r}(p^{\prime})\right]=\pi_{r,r}(p_{r^{*}})$. The price at which $\pi_{r,s}(p)$ is equal to $\pi_{r,r}(p_{r^{*}})$ was defined in Section 2.2 to be $p^{\dagger}$, which is the solution to the following quadratic equation:
(58) $\pi_{r,r}(p_{r^{*}})$ $=\pi_{r,s}(p^{\dagger})$
$\frac{(1-c_{r})^{2}}{4}$ $=\alpha p^{\dagger}(1-p^{\dagger})$
$(p^{\dagger})^{2}-p^{\dagger}+\frac{(1-c_{r})^{2}}{4\alpha}$ $=0$
$p^{\dagger}$ $=\frac{1+\sqrt{1-\frac{(1-c_{r})^{2}}{\alpha}}}{2}$
Where we choose the larger solution, as the smaller solution satisfies the following inequalities
(59) $\frac{1-\sqrt{1-\frac{(1-c_{r})^{2}}{\alpha}}}{2}<\frac{1}{2}<\frac{1+c_{r}}{2}=p_{r^{*}}$
and therefore drops when we consider that we are handling the case of $p_{r^{*}}<p$. Note that, in order for $p^{\dagger}$ to be well defined, we need the discriminant to be positive which occurs precisely when $\alpha\geq(1-c_{r})^{2}$; we will return to this later in this subsection. Thus for an equilibrium in this case, we need that $p\in\left[\max\{0,p_{r^{*}},p_{s_{\text{ind}}}\},\min\{1,p^{\dagger},p_{s^{*}}\}\right]$. Note that $p^{\dagger}\leq 1$ and $p_{r^{*}},p_{s_{\text{ind}}}\geq 0$ always, so we can drop those conditions.
From (46), we have that $p_{s_{\text{ind}}}\geq p_{r^{*}}\iff\alpha\geq\alpha_{r^{*}}$. We won’t explicitly simplify the condition $p_{s^{*}}\leq p^{\dagger}$ here; we delay that until (76). For now we will simply remark that $\pi_{r,s}(p)$ is maximized at $p=\frac{1}{2}$, which follows because
$\frac{\partial\pi_{r,s}}{\partial p}$ $=\frac{\partial}{\partial p}\left(\alpha p(1-p)\right)$
$=\alpha(1-2p)\stackrel{{\scriptstyle\text{set}}}{{=}}0\iff p=\frac{1}{2}$
So, because $\frac{1}{2}\leq p_{s^{*}},p^{\dagger}$ and $\pi_{r,s}$ is a concave quadratic, we have that
(60) $p_{s^{*}}\leq p^{\dagger}\iff\pi_{r,r}(p_{r^{*}})=\pi_{r,s}(p^{\dagger})\leq\pi_{r,s}(p_{s^{*}})$
We’ll check to make sure that the interval of prices is nonempty in each of the possible cases:
Interval is $[p_{s_{\text{ind}}},\min\left\{p_{s^{*}},p^{\dagger}\right\}]$.
Per (47), we have that $p_{s_{\text{ind}}}\leq p_{s^{*}}\iff\alpha\leq 1-c_{s}$. However, we have not found the conditions under which $p_{s_{\text{ind}}}\leq p^{\dagger}$. By the same argument as in (60), we have the following equivalences:
$p_{s_{\text{ind}}}$ $\leq p^{\dagger}$
$\pi_{r,r}(p_{r^{*}})=\pi_{r,s}(p^{\dagger})$ $\leq\pi_{r,s}(p_{s_{\text{ind}}})$
(61) $\left(\frac{1-c_{r}}{2}\right)^{2}$ $\leq\frac{\alpha c_{s}}{1-\alpha}\left(1-\frac{c_{s}}{1-\alpha}\right)$
Now, recall that
$p_{s_{\text{ind}}}\geq p_{r^{*}}\iff\alpha\geq\alpha_{r^{*}}\geq\alpha^{\dagger}_{r}=\frac{1-c_{r}}{1+c_{r}}\geq(1-c_{r})^{2}$
which implies that $p^{\dagger}$ is well defined and that $\pi_{r,s}(\frac{1}{2})\geq\pi_{r,r}(p_{r^{*}})$. We see upon substituting the endpoint $\alpha=\alpha^{\dagger}_{r}=1-\frac{2c_{s}}{1+c_{r}}$ into (61) that
$\left(\frac{1-c_{r}}{2}\right)^{2}$ $\leq\frac{\alpha^{\dagger}_{r}c_{s}}{1-\alpha^{\dagger}_{r}}\left(1-\frac{c_{s}}{1-\alpha^{\dagger}_{r}}\right)$
$\left(\frac{1-c_{r}}{2}\right)^{2}$ $\leq\left(\frac{1-2c_{s}+c_{r}}{1+c_{r}}\right)\frac{c_{s}}{\frac{2c_{s}}{1+c_{r}}}\left(1-\frac{c_{s}}{\frac{2c_{s}}{1+c_{r}}}\right)$
$\left(\frac{1-c_{r}}{2}\right)^{2}$ $\leq\left(\frac{1-c_{r}}{2}\right)\left(\frac{1-2c_{s}+c_{r}}{2}\right)$
$1-c_{r}$ $\leq 1+c_{r}-2c_{s}$
$c_{s}$ $\leq c_{r}$
Furthermore, we show in (85) and (86) that $\pi_{r,s}(p_{s_{\text{ind}}},\alpha)$ has a unique maximum at $\alpha=\frac{1-c_{s}}{1+c_{s}}<\alpha_{r^{*}}$, which is sufficient to imply that there exists some $\alpha^{\dagger}_{s}\in\left[\alpha_{r^{*}},1-c_{s}\right]$ such that $\alpha\leq\alpha^{\dagger}_{s}\iff p_{s_{\text{ind}}}\leq p^{\dagger}$. For completeness, explicitly solving (61) yields the closed-form
(62) $\alpha^{\dagger}_{s}\equiv\frac{c_{s}(1-c_{s})+\frac{(1-c_{r})^{2}}{2}+c_{s}\sqrt{(1-c_{s})^{2}-(1-c_{r})^{2}}}{2c_{s}+\frac{(1-c_{r})^{2}}{2}}\in\left[\alpha_{r^{*}},1-c_{s}\right]$
Therefore, the interval is feasible as long as $\alpha\in\left[\alpha_{r^{*}},\alpha^{\dagger}_{s}\right]$.
Interval is $[p_{r^{*}},p_{s^{*}}]$.
Per (52), we have that $p_{r^{*}}\leq p_{s^{*}}\iff\alpha\geq\alpha_{s^{*},r^{*}}$, so $[p_{r^{*}},p_{s^{*}}]$ is nonempty whenever $\alpha\in\left[\alpha_{s^{*},r^{*}},\alpha_{r^{*}}\right]$. Note we showed in (53) that $\alpha_{s^{*},r^{*}}\leq\alpha_{r^{*}}$ always.
Interval is $[p_{r^{*}},p^{\dagger}]$.
Interestingly, we find the following condition for $p_{r^{*}}\leq p^{\dagger}$:
$p^{\dagger}$ $\geq p_{r^{*}}$
$\frac{1+\sqrt{1-\frac{(1-c_{r})^{2}}{\alpha}}}{2}$ $\geq\frac{1+c_{r}}{2}$
$\sqrt{1-\frac{(1-c_{r})^{2}}{\alpha}}$ $\geq c_{r}$
$\frac{(1-c_{r})^{2}}{\alpha}$ $\leq 1-c_{r}^{2}$
(63) $\alpha$ $\geq\frac{1-c_{r}}{1+c_{r}}=\alpha^{\dagger}_{r}$
Thus, $[p_{r^{*}},p^{\dagger}]$ is nonempty so long as $\alpha\in\left[\alpha^{\dagger}_{r},\alpha_{r^{*}}\right]$. Note we showed in (51) that $\alpha^{\dagger}_{r}\leq\alpha_{r^{*}}$ always.
Reconciling equilibria with case (i) where $c_{s}\leq c_{s^{*}}$.
We’ve so far shown that $p\in[p_{r^{*}},p_{s^{*}}]$ or $p\in[p_{r^{*}},p^{\dagger}]$ could be an equilibrium if $\alpha\in\left[\min\left\{\alpha_{s^{*},r^{*}},\alpha^{\dagger}_{r}\right\},1-c_{s}\right]$. In (56) and (57), we subdivided the equilibria into two cases, when $c_{s}\leq c_{s^{*}}$ and $c_{s}\geq c_{s^{*}}$. Though we have not explicitly simplified the condition $p_{s^{*}}\leq p^{\dagger}$, it would be prudent to consider what equilibria exist when $\alpha$ is between $\alpha^{\dagger}_{r}$ and $\alpha_{s^{*},r^{*}}$.
More specifically, suppose $c_{s}\leq c_{s^{*}}$. We will show here the following claim: $p_{s^{*}}\leq p^{\dagger}$ for all $\alpha\in\left[\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\right]$. Recalling our condition from (60), we have that
$p_{s^{*}}\leq p^{\dagger}\iff\pi_{r,r}(p_{r^{*}})=\frac{1}{4}(1-c_{r})^{2}\leq\frac{\alpha}{4}\left(1-\left(\frac{c_{s}}{1-\alpha}\right)^{2}\right)=\pi_{r,s}(p_{s^{*}},\alpha)$
In this case, we have
$\pi_{r,s}\left(p_{s^{*}},\alpha_{s^{*},r^{*}}\right)$ $=\frac{\alpha_{s^{*},r^{*}}}{4}\left(1-\left(\frac{c_{s}}{1-\alpha_{s^{*},r^{*}}}\right)^{2}\right)$
$=\frac{1}{4}\left(\frac{c_{r}-c_{s}}{c_{r}}\right)\left(1-\left(\frac{c_{s}}{\frac{c_{s}}{c_{r}}}\right)^{2}\right)$
$=\frac{1}{4}\left(\frac{c_{r}-c_{s}}{c_{r}}\right)\left(1-c_{r}^{2}\right)$
$\geq\frac{1}{4}\left(\frac{1-c_{r}}{1+c_{r}}\right)(1-c_{r})(1+c_{r})$
$=\frac{(1-c_{r})^{2}}{4}=\pi_{r,r}(p_{r^{*}})$
Similarly, we can see that
$\pi_{r,s}\left(p_{s^{*}},\alpha^{\dagger}_{r}\right)$ $=\frac{\alpha^{\dagger}_{r}}{4}\left(1-\left(\frac{c_{s}}{1-\alpha^{\dagger}_{r}}\right)^{2}\right)$
$=\frac{1}{4}\left(\frac{1-c_{r}}{1+c_{r}}\right)\left(1-\left(\frac{c_{s}}{1-\frac{1-c_{r}}{1+c_{r}}}\right)^{2}\right)$
$\geq\frac{1}{4}\left(\frac{1-c_{r}}{1+c_{r}}\right)\left(1-\left(\frac{c_{s}}{1-\frac{c_{r}-c_{s}}{c_{r}}}\right)^{2}\right)$
$=\frac{(1-c_{r})^{2}}{4}=\pi_{r,r}(p_{r^{*}})$
Thus, to show the claim, all we need to show is that $\pi_{r,s}(p_{s^{*}},\alpha)$ is lower bounded by its value at the endpoints for $\alpha\in\left[\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\right]$. We do this by finding the derivative:
$\frac{\partial\pi_{r,s}(p_{s^{*}},\alpha)}{\partial\alpha}$ $=\frac{\partial}{\partial\alpha}\left(\frac{\alpha}{4}\left(1-\left(\frac{c_{s}}{1-\alpha}\right)^{2}\right)\right)$
$=\frac{1}{4}\left(1-\left(\frac{c_{s}}{1-\alpha}\right)^{2}\right)-\frac{\alpha}{4}\left(\frac{2c_{s}^{2}}{(1-\alpha)^{3}}\right)$
$=\frac{1}{4}\left[1-\frac{c_{s}^{2}(1-\alpha)+2\alpha c_{s}^{2}}{(1-\alpha)^{3}}\right]$
(64) $=\frac{1}{4}\left[1-\frac{c_{s}^{2}(1+\alpha)}{(1-\alpha)^{3}}\right]$
Note that the $\tilde{\alpha}$ that maximizes $\pi_{r,s}(p_{s^{*}},\alpha)$ (which will appear in the future in Lemma 4 where we show that $\overline{\alpha}=\tilde{\alpha}$) satisfies the following –
$c_{s}^{2}=\frac{(1-\tilde{\alpha})^{3}}{1+\tilde{\alpha}}$
Because the function $f(x)=\frac{(1-x)^{3}}{1+x}$ is a bijection from (0,1][0,1)0101(0,1$\to$[0,1)( 0 , 1 ] → [ 0 , 1 ), this importantly implies that $\pi_{r,s}(p_{s^{*}},\alpha)$ has a unique maximum for $c_{s}^{2}\in[0,1)$. Now, we claim that $\pi_{r,s}$ is decreasing at $\alpha_{s^{*},r^{*}}$, which follows because –
00 0 $\leq\frac{\partial\pi_{r,s}(p_{s^{*}},\alpha)}{\partial\alpha}\biggr{|}_{\alpha=\alpha_{s^{*},r^{*}}}=\frac{1}{4}\left[1-\frac{c_{s}^{2}\left(\frac{2c_{r}-c_{s}}{c_{r}}\right)}{\left(\frac{c_{s}}{c_{r}}\right)^{3}}\right]$
00 0 $\leq 1-\frac{c_{r}^{2}(2c_{r}-c_{s})}{c_{s}}$
$\frac{c_{s}}{c_{r}^{2}}$ $\leq 2c_{r}-c_{s}$
$\alpha_{s^{*}}=1-2c_{r}+c_{s}$ $\leq\frac{c_{r}-c_{s}}{c_{r}^{2}}=\frac{\alpha_{s^{*},r^{*}}}{c_{r}}$
Which always holds because, as we showed in (54), when $c_{s}\leq c_{s^{*}}$ we have $\alpha_{s^{*}}\leq\alpha_{s^{*},r^{*}}\leq\frac{\alpha_{s^{*},r^{*}}}{c_{r}}$ always. This is sufficient to conclude the proof of the claim. Given the claim, we can now conclude that when $\alpha\in\left[\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\right]$, the region of equilibria we have found is $p\in\left[p_{s_{\text{ind}}},p_{s^{*}}\right]$, which is the same as we found before. Note that $p^{\dagger}$ is well defined for all $\alpha\geq(1-c_{r})^{2}$, which is implied by $\alpha\geq\alpha^{\dagger}_{r}$:
(65) $\alpha^{\dagger}_{r}=\frac{1-c_{r}}{1+c_{r}}$ $\geq(1-c_{r})^{2}$
11 11 $\geq(1-c_{r})(1+c_{r})$
$c_{r}^{2}$ $\geq 0$
Reconciling equilibria with case (ii) where $c_{s}\geq c_{s^{*}}$.
Switching the direction of the inequalities in the proof from the previous subsection yields the analogous result that if $c_{s}\geq c_{s^{*}}$, $p^{\dagger}\leq p_{s^{*}}$ for all $\alpha\in\left[\alpha_{s^{*},r^{*}},\alpha^{\dagger}_{r}\right]$. We previously found that for all $\alpha\leq\alpha^{\dagger}_{r}$, the equilibria are $p\in\left[p_{s_{\text{ind}}},p_{r_{\text{ind}}}\right]$. However, by the same argument as in (60) we have that
$\frac{1}{2}\leq p^{\dagger}\leq p_{r_{\text{ind}}}\iff\pi_{r,s}(p^{\dagger})\geq\pi_{r,s}(p_{r_{\text{ind}}})\iff\pi_{r,r}(p_{r^{*}})\geq\pi_{r,r}(p_{r_{\text{ind}}})$
which always holds, thus we have found no new equilibria in this region of parameter space.
Conclusion and summary of equilibria.
Combining with the equilibria from the first case from (56) and (57), we have the following shared-price equilibria where the independent seller fulfills –
Case (i), $c_{s}\leq c_{s^{*}}$:
(66) $p\in\begin{cases}\left[p_{s_{\text{ind}}},p_{r_{\text{ind}}}\right]&\alpha\leq\alpha_{s^{*}}\\ \left[p_{s_{\text{ind}}},p_{s^{*}}\right]&\alpha_{s^{*}}\leq\alpha\leq\alpha_{s^{*},r^{*}}\\ \left[p_{s_{\text{ind}}},\min\left\{p_{s^{*}},p^{\dagger}\right\}\right]&\alpha_{s^{*},r^{*}}\leq\alpha\leq\alpha^{\dagger}_{s}\\ \end{cases}$
Case (ii), $c_{s}\geq c_{s^{*}}$:
(67) $p\in\begin{cases}\left[p_{s_{\text{ind}}},p_{r_{\text{ind}}}\right]&\alpha\leq\alpha^{\dagger}_{r}\\ \left[p_{s_{\text{ind}}},\min\left\{p_{s^{*}},p^{\dagger}\right\}\right]&\alpha^{\dagger}_{r}\leq\alpha\leq\alpha^{\dagger}_{s}\\ \end{cases}$
D.3. Equilibria where $p_{r}<p_{s}$
Condition IV is always satisfied; because for all $\varepsilon>0$, $p_{s}+\varepsilon>p_{s}>p_{r}$, the independent seller achieves a payoff of 00 both at price $p_{s}$ and price $p_{s}+\varepsilon$. Together, conditions I and II imply that the retailer must set price $p_{r^{*}}$ so as to maximize their payoff function. Otherwise, if they are not at the optimal price, they would be able to increase their payoff by changing their price.
Furthermore, condition II requires that it is not the retailer would not achieve a higher payoff by deviating to some price $p_{r}>p_{s}$ so as to let the independent seller fulfill demand in equilibrium. Formally, we need that
$\pi_{r,s}(p_{s})\leq\pi_{r,r}(p_{r^{*}})$
However, this is exactly the equation we solved in (58), so we have that $p_{s}$ must be greater than $p^{\dagger}$ (recall from (59) that the smaller root is always less than $p_{r^{*}}$). Note that we also must have that $p_{s}\geq p_{r^{*}}$, and in (63) we showed that $p_{r^{*}}\leq p^{\dagger}$ is equivalent to $\alpha\geq\alpha^{\dagger}_{r}$.
We move on to condition III now. Recall that for any $p_{s}>p_{r}$ we have that $\pi_{s}(p_{r},p_{s})=0$. Thus, condition III simplifies to
$\max_{p_{s}^{\prime}\in[0,p_{r^{*}}]}\left[\pi_{s}(p_{r^{*}},p_{s}^{\prime})\right]\leq 0$
The maximum of the independent seller’s payoff on the interval $[0,p_{r^{*}}]$ is attained at
$\arg\max_{p_{s}^{\prime}\in[0,p_{r^{*}}]}\left[\pi_{s}(p_{r^{*}},p_{s}^{\prime})\right]=\begin{cases}p_{r^{*}}&p_{r^{*}}\leq p_{s^{*}}\\ p_{s^{*}}&p_{r^{*}}>p_{s^{*}}\end{cases}$
because $\pi_{s}(\cdot,p_{s}^{\prime})$ is maximized at $p_{s^{*}}$. Consider the second case when $p_{r^{*}}>p_{s^{*}}$. Recall from (9) that
$\pi_{s,s}(p_{s^{*}})=\begin{cases}\frac{1-\alpha}{4}\left(1-\frac{c_{s}}{1-\alpha}\right)^{2}&\alpha<1-c_{s}\\ 0&\alpha\geq 1-c_{s}\end{cases}$
Thus, we find immediately that when $p_{r^{*}}>p_{s^{*}}$, condition III never holds if $\alpha<1-c_{s}$ and always holds if $\alpha\geq 1-c_{s}$. However, if $\alpha\geq 1-c_{s}$, we have that $p_{s^{*}}\geq 1>p_{r^{*}}$, which is a contradiction. Thus there are no equilibria where $p_{r^{*}}>p_{s^{*}}$.
On the other hand, $\arg\max_{p_{s}^{\prime}\in[0,p_{r^{*}}]}\left[\pi_{s}(p_{r^{*}},p_{s}^{\prime})\right]=p_{r^{*}}$ if $p_{r^{*}}\leq p_{s^{*}}$, which we showed in (52) is equivalent to $\alpha\geq\alpha_{s^{*},r^{*}}$. With this, condition III is satisfied in this case if
$\pi_{s,s}(p_{r^{*}})$ $\leq 0$
$\left((1-\alpha)\left(\frac{1+c_{r}}{2}\right)-c_{s}\right)\left(1-\frac{1+c_{r}}{2}\right)$ $\leq 0$
$\left(\frac{(1-\alpha)(1+c_{r})-2c_{s}}{2}\right)\left(\frac{1-c_{r}}{2}\right)$ $\leq 0$
$(1-\alpha)(1+c_{r})$ $\leq 2c_{s}$
(68) $\frac{1+c_{r}}{2}$ $\leq\frac{c_{s}}{1-\alpha}$
(69) $\alpha$ $\geq 1-\frac{2c_{s}}{1+c_{r}}=\alpha_{r^{*}}$
As a remark, see that (68) can be rewritten as $p_{r^{*}}\leq p_{s_{\text{ind}}}$, which is obvious in hindsight because $p_{s_{\text{ind}}}$ is by construction the price below which the independent seller makes nonpositive payoff. We showed in (51) that $\alpha^{\dagger}_{r}\leq\alpha_{r^{*}}$, so $\alpha\geq\alpha_{r^{*}}$ is sufficient to imply that $p_{r^{*}}\leq p^{\dagger}$ per (63). Additionally, recall from (53) that $\alpha_{s^{*},r^{*}}\leq\alpha_{r^{*}}$.
In summary, we have a continuum of equilibria $(p_{r^{*}},p_{s}\geq p^{\dagger})$ whenever $\alpha\geq\alpha_{r^{*}}$.
D.4. Equilibria where $p_{s}<p_{r}$
In this case, condition II is always satisfied; the retailer does not fulfill any demand either way as, for all $\varepsilon>0$, $p_{r}+\varepsilon>p_{r}>p_{s}$. As in the previous case, conditions III and IV together imply that the independent seller must set price $p_{s^{*}}$. Condition IV requires that the independent seller does not prefer to let the retailer fulfill demand in equilibrium, in which case they would earn a payoff of 00. To satisfy this, we need only that the independent seller sets a price $p_{s}\geq c_{s}$, which always holds for price $p_{s^{*}}$.
Thus, all that remans to check is condition I. Simplifying the condition, we need that
$\max_{p_{r}^{\prime}\in[0,p_{s^{*}}]}\left[\pi_{r,r}(p_{r}^{\prime})\right]\leq\pi_{r,s}(p_{s^{*}})$
Once again, there are two cases for the choice of $p_{r}^{\prime}$ that maximizes the retailer’s payoff function on the interval $[0,p_{s^{*}}]$
$\arg\max_{p_{r}^{\prime}\in[0,p_{s^{*}}]}\left[\pi_{r}(p_{r}^{\prime},p_{s^{*}})\right]=\begin{cases}p_{s^{*}}&p_{s^{*}}\leq p_{r^{*}}\\ p_{r^{*}}&p_{s^{*}}>p_{r^{*}}\end{cases}$
Consider the second case where $\arg\max_{p_{r}^{\prime}\in[0,p_{s^{*}}]}\left[\pi_{r}(p_{r}^{\prime},p_{s^{*}})\right]=p_{r^{*}}$. From (52), recall that $p_{s^{*}}>p_{r^{*}}\iff\alpha>\alpha_{s^{*},r^{*}}$. Recalling our analysis from (60), condition I then simplifies to
$\pi_{r,r}(p_{r^{*}})$ $\leq\pi_{r,s}(p_{s^{*}})$
$\pi_{r,s}(p^{\dagger})$ $\leq\pi_{r,s}(p_{s^{*}})$
$p_{s^{*}}$ $\leq p^{\dagger}$
Note that $\alpha>\alpha^{\dagger}_{s}\implies p^{\dagger}<p_{s_{\text{ind}}}<p_{s^{*}}$ by construction of $\alpha^{\dagger}_{s}$, thus $\alpha\leq\alpha^{\dagger}_{s}$ is necessary for condition I to hold. Thus, we have a continuum of equilibria $(p_{r}>p_{s^{*}},p_{s^{*}})$ whenever $\alpha\in\left[\alpha_{s^{*},r^{*}},\alpha^{\dagger}_{s}\right]$ and $p_{s^{*}}\leq p^{\dagger}$ holds. Additionally, recall that if $c_{s}\geq c_{s^{*}}$, we showed above that $p_{s^{*}}\leq p^{\dagger}$ can only hold when $\alpha\geq\alpha^{\dagger}_{r}$.
Now consider the first case where $\arg\max_{p_{r}^{\prime}\in[0,p_{s^{*}}]}\left[\pi_{r}(p_{r}^{\prime},p_{s^{*}})\right]=p_{s^{*}}$. From (52), recall that $p_{s^{*}}\leq p_{r^{*}}\iff\alpha\leq\alpha_{s^{*},r^{*}}$. On the other hand, we have that condition I becomes
$\pi_{r,r}(p_{s^{*}})$ $\leq\pi_{r,s}(p_{s^{*}})$
$(1-p_{s^{*}})(p_{s^{*}}-c_{r})$ $\leq\alpha p_{s^{*}}(1-p_{s^{*}})$
$(1-\alpha)p_{s^{*}}$ $\leq c_{r}$
$\frac{1-\alpha+c_{s}}{2}$ $\leq c_{r}$
$\alpha$ $\geq 1-2c_{r}+c_{s}=\alpha_{s^{*}}$
Thus, we have a continuum of equilibria $(p_{r}>p_{s^{*}},p_{s^{*}})$ whenever $\alpha\in\left[\alpha_{s^{*}},\alpha_{s^{*},r^{*}}\right]$. Note that it’s not always the case that $\alpha_{s^{*}}\leq\alpha_{s^{*},r^{*}}$, but when this holds, such equilibria exist. Additionally, it’s clear that $\alpha_{s^{*},r^{*}}=1-\frac{c_{s}}{c_{r}}<1-c_{s}$ always.
Appendix E Refining and interpreting equilibria - proofs
E.1. Admissibility
For the independent seller, playing $p_{s}<p_{s_{\text{ind}}}$ is weakly dominated by playing a price $p_{s}=1$, because their payoffs are equal if the retailer sets a price $p_{r}\leq p_{s}$ but for $p_{r}\in(p_{s},1)$ we have
$\pi_{s,s}(p_{s})=\pi_{s}(p_{r},p_{s})\leq\pi_{s}(p_{r},1)=0$
because they achieve negative payoff when fulfilling demand below their breakeven price. Playing a price $p_{s}>p_{s^{*}}$ is dominated by playing $p_{s^{*}}$ because, for all $p_{r}>p_{s}$,
$\pi_{s,s}(p_{s})=\pi_{s}(p_{r},p_{s})\leq\pi_{s}(p_{r},p_{s^{*}})=\pi_{s,s}(p_{s^{*}})$
where the inequality holds because $p_{s}>p_{s^{*}}$. For any price $p_{r}\leq p_{s}$, we have
$0=\pi_{s,r}(p_{s})=\pi_{s}(p_{r},p_{s})\leq\pi_{s}(p_{r},p_{s^{*}})$
which implies weak dominance as desired. However, consider for instance two prices $p_{s_{\text{ind}}}<p_{s_{1}}<p_{s_{2}}<p_{s^{*}}$, we claim neither dominates the other. If $p_{r}>p_{s_{2}}$,
$\pi_{s,s}(p_{s_{1}})=\pi_{s}(p_{r},p_{s_{1}})<\pi_{s}(p_{r},p_{s_{2}})=\pi_{s,s}(p_{s_{2}})$
because $p_{s_{1}}<p_{s_{2}}<p_{s^{*}}$. On the other hand if $p_{r}\in(p_{s_{1}},p_{s_{2}})$,
$\pi_{s,s}(p_{s_{1}})=\pi_{s}(p_{r},p_{s_{1}})>\pi_{s}(p_{r},p_{s_{2}})=0$
Thus, in any region of parameter space, the only admissible strategies for the independent seller are $p_{s}\in(p_{s_{\text{ind}}},p_{s^{*}})$, except for a technical edge case when $p_{s_{\text{ind}}}\geq p_{s^{*}}\geq 1$ which we will address later.
A similar result holds for the retailer, except it’s not always the case that $p_{r_{\text{ind}}}\leq p_{r^{*}}$ (which occurs precisely when $\alpha\leq\alpha^{\dagger}_{r}$ per (49)). We claim that $p_{r}=\min\left\{p_{r_{\text{ind}}},p_{r^{*}}\right\}$ dominates all $p_{r}^{\prime}<p_{r}$. Indeed, for $p_{s}\leq p_{r}^{\prime}$, the retailer’s payoff is the same. For $p_{s}\geq p_{r}$, the retailer’s payoff is
$\pi_{r,r}(p_{r}^{\prime})=\pi_{r}(p_{r}^{\prime},p_{s})\leq\pi_{r}(p_{r},p_{s})=\pi_{r,r}(p_{r})$
where the inequality holds because $p_{r}^{\prime}<p_{r}\leq p_{r^{*}}$. For $p_{s}\in(p_{r}^{\prime},p_{r})$, we have
$\pi_{r,r}(p_{r}^{\prime})=\pi_{r}(p_{r}^{\prime},p_{s})\leq\pi_{r}(p_{r},p_{s})=\pi_{r,s}(p_{s})$
where the inequality holds because $p_{r}^{\prime}<p_{s}<p_{r}\leq p_{r_{\text{ind}}}$.
Similarly, $p_{r}=\max\left\{p_{r_{\text{ind}}},p_{r^{*}}\right\}$ dominates all $p_{r}^{\prime}>p_{r}$. Indeed, for $p_{s}\leq p_{r}$, the retailer’s payoff is the same. For $p_{s}\geq p_{r}^{\prime}$, the retailer’s payoff is
$\pi_{r,r}(p_{r}^{\prime})=\pi_{r}(p_{r}^{\prime},p_{s})\leq\pi_{r}(p_{r},p_{s})=\pi_{r,r}(p_{r})$
where the inequality holds because $p_{r}^{\prime}>p_{r}\geq p_{r^{*}}$. For $p_{s}\in(p_{r},p_{r}^{\prime})$, we have
$\pi_{r,s}(p_{s})=\pi_{r}(p_{r}^{\prime},p_{s})\leq\pi_{r}(p_{r},p_{s})=\pi_{r,r}(p_{r})$
where the inequality holds because $p_{r}^{\prime}>p_{s}>p_{r}\geq p_{r_{\text{ind}}}$. However, we can show fairly easily that for any $\min\left\{p_{r_{\text{ind}}},p_{r^{*}}\right\}<p_{r_{1}}<p_{r_{2}}<\max\left\{p_{r_{\text{ind}}},p_{r^{*}}\right\}$, neither weakly dominates the other. Suppose $p_{r_{\text{ind}}}<p_{r^{*}}$, then for $p_{s}>p_{r_{2}}$ we have
$\pi_{r}(p_{r_{1}},p_{s})=\pi_{r,r}(p_{r_{1}})\leq\pi_{r}(p_{r_{2}},p_{s})=\pi_{r,r}(p_{r_{2}})$
because $p_{r_{1}}<p_{r_{2}}<p_{r^{*}}$. While, for, $p_{s}\in(p_{r_{1}},p_{r_{2}})$, we have
$\pi_{r,s}(p_{r_{2}})=\pi_{r}(p_{r_{2}},p_{s})\leq\pi_{r}(p_{r_{1}},p_{s})=\pi_{r,r}(p_{r_{1}})$
because $p_{r_{\text{ind}}}<p_{r_{1}}<p_{r_{2}}$. Reversing the directions of the inequalities proves that there are no dominant strategies for any $p_{r^{*}}<p_{r_{1}}<p_{r_{2}}<p_{r_{\text{ind}}}$.
With this in mind, we will step through each of the Nash equilibria that we found in (66) and (67) and remove those strategies that are weakly dominated.
First consider an equilibrium in the regime where $\alpha\leq\min\left\{\alpha_{s^{*}},\alpha^{\dagger}_{r}\right\}$. Here, the Nash equilibrium is $p\in[p_{s_{\text{ind}}},p_{r_{\text{ind}}}]$. Because the retailer’s admissible strategies fall in the interval $p_{r}\in[p_{r_{\text{ind}}},p_{r^{*}}]$, the only admissible equilibrium price is $p=p_{r_{\text{ind}}}$.
Now, suppose that $\alpha\in\left[\alpha_{s^{*}},\alpha_{s^{*},r^{*}}\right]$, in which case the Nash equilibria are $(p_{r}\geq p_{s^{*}},p_{s}=p_{s^{*}})\cup(p\in[p_{s_{\text{ind}}},p_{s^{*}}])$. We showed that in this region of parameter space, we have $p_{s_{\text{ind}}}\leq p_{s^{*}}\leq p_{r^{*}},p_{r_{\text{ind}}}$. Notably, this implies that the entire continuum of shared-price equilibria $p\in[p_{s_{\text{ind}}},p_{s^{*}}]$ are inadmissible because it is entirely disjoint from the admissible region for the retailer, $p_{r}\in\left[\min\left\{p_{r_{\text{ind}}},p_{r^{*}}\right\},\max\left\{p_{r_{\text{ind}}},p_{r^{*}}\right\}\right]$. Thus, the admissible equilibria are $(p_{r}\in\left[p_{r_{\text{ind}}},p_{r^{*}}\right],p_{s}=p_{s^{*}})$ if $\alpha\in\left[\alpha_{s^{*}},\alpha^{\dagger}_{r}\right]$ and $(p_{r}\in\left[p_{r^{*}},p_{r_{\text{ind}}}\right],p_{s}=p_{s^{*}})$ if $\alpha\in\left[\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\right]$.
Now, suppose $p_{s^{*}}\leq p^{\dagger}$ and $\alpha\in\left[\max\left\{\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\right\},\alpha_{r^{*}}\right]$. As before, the Nash equlibria are $(p_{r}\geq p_{s^{*}},p_{s}=p_{s^{*}})\cup(p\in\left[p_{s_{\text{ind}}},p_{s^{*}}\right])$. However, this time, we have $p_{s_{\text{ind}}}\leq p_{r^{*}}\leq p_{s^{*}}\leq p^{\dagger}\leq p_{r_{\text{ind}}}$. Here, the retailer’s admissible prices are $p_{r}\in[p_{r^{*}},p_{r_{\text{ind}}}]$, so the remaining admissible equilibria are $(p_{r}\in[p_{s^{*}},p_{r_{\text{ind}}}],p_{s}=p_{s^{*}})\cup(p\in\left[p_{r^{*}},p_{s^{*}}\right])$.
The arguments of the previous paragraph did not depend on the relative ordering of $p^{\dagger}$ and $p_{s^{*}}$, and we know that $p_{r_{\text{ind}}}\geq p^{\dagger}\geq p_{r^{*}}$ per (63) because $\alpha\geq\alpha^{\dagger}_{r}$. Thus the admissible equilibria when $\alpha\in\left[\max\left\{\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\right\},\alpha_{r^{*}}\right]$ and $p^{\dagger}\leq p_{s^{*}}$ are simply the shared prices $p\in\left[p_{r^{*}},p^{\dagger}\right]$.
Consider the case when $\alpha\in\left[\alpha_{r^{*}},\alpha^{\dagger}_{s}\right]$, where the equilibria are $(p_{r}=p_{r^{*}},p_{s}\geq p^{\dagger})\cup(p\in[p_{s_{\text{ind}}},\min\{p^{\dagger},p_{s^{*}}\}])$. In this case, $p_{r^{*}}\leq p_{s_{\text{ind}}}\leq p_{s^{*}},p^{\dagger}\leq p_{r_{\text{ind}}}$ always. If $p_{s^{*}}\leq p^{\dagger}$ also, then the only admissible equilibria are $p\in[p_{s_{\text{ind}}},p_{s^{*}}]$ as prices $p_{s}\geq p_{s^{*}}$ are inadmissible. On the other hand, if $p^{\dagger}\leq p_{s^{*}}$, the admissible equilibria are $(p_{r}=p_{r^{*}},p_{s}\in[p^{\dagger},p_{s^{*}}])\cup(p\in[p_{s_{\text{ind}}},p^{\dagger}])$.
Now suppose that $\alpha\in[\alpha^{\dagger}_{s},1-c_{s}]$, here the only Nash equilibrium is $(p_{r}=p_{r^{*}},p_{s}\geq p^{\dagger})$. In this region of parameter space, we always have that $p^{\dagger}\leq p_{s_{\text{ind}}}\leq p_{s^{*}}\leq 1$ by construction of $\alpha^{\dagger}_{s}$, so the admissible Nash equilibria are $(p_{r}=p_{r^{*}},p_{s}\in[p_{s_{\text{ind}}},p_{s^{*}}])$.
However, when $\alpha\geq 1-c_{s}\geq\alpha^{\dagger}_{s}$, we have that $p_{s_{\text{ind}}}\geq p_{s^{*}}\geq 1$ and the Nash equilibrium is again given by $(p_{r}=p_{r^{*}},p_{s}\geq p^{\dagger})$. In this limit, setting $p_{s}=1$ weakly dominates setting any $p_{s}^{\prime}<1$ because $p_{s_{\text{ind}}}\geq 1>p_{s}^{\prime}$ which implies that the independent seller would achieve negative payoff for each unit sold at price $p_{s}^{\prime}$ and would prefer to achieve 0 payoff by setting $p_{s}=1$. Thus, the only admissible equilibrium when $\alpha\geq 1-c_{s}$ is $(p_{r}=p_{r^{*}},p_{s}=1)$. Note that this edge case is technical and unimportant, as no matter what price the independent seller sets, the result remains the same – when $\alpha$ is high, there is an equilibrium where the retailer fulfills demand.
Summarizing, we have narrowed our focus to the following admissible equilibria:
Shared equilibria, irrespective of ordering of $c_{s}$ and $c_{s^{*}}$:
(70) $(p_{r},p_{s})\in\begin{cases}(p_{s},[p_{s_{\text{ind}}},p^{\dagger}])\cup(p_{r^{*}},[p^{\dagger},p_{s^{*}}])&\left[\alpha_{r^{*}}\leq\alpha\leq\alpha^{\dagger}_{s}\right]\land[p^{\dagger}\leq p_{s^{*}}]\\ (p_{s},[p_{s_{\text{ind}}},p_{s^{*}}])&\left[\alpha_{r^{*}}\leq\alpha\leq\alpha^{\dagger}_{s}\right]\land[p_{s^{*}}\leq p^{\dagger}]\\ (p_{r^{*}},[p_{s_{\text{ind}}},p_{s^{*}}])&\alpha^{\dagger}_{s}\leq\alpha\leq 1-c_{s}\\ (p_{r^{*}},1)&1-c_{s}\leq\alpha\\ \end{cases}$
Case (i), $c_{s}\leq c_{s^{*}}$:
(71) $(p_{r},p_{s})\in\begin{cases}(p_{r_{\text{ind}}},p_{r_{\text{ind}}})&\alpha\leq 1-2c_{r}+c_{s}\\ ([p_{r_{\text{ind}}},p_{r^{*}}],p_{s^{*}})&\alpha_{s^{*}}\leq\alpha\leq\alpha^{\dagger}_{r}\\ ([p_{r^{*}},p_{r_{\text{ind}}}],p_{s^{*}})&\alpha^{\dagger}_{r}\leq\alpha\leq\alpha_{s^{*},r^{*}}\\ (p_{s},\left[p_{r^{*}},p_{s^{*}}\right])\cup([p_{s^{*}},p_{r_{\text{ind}}}],p_{s^{*}})&\left[\alpha_{s^{*},r^{*}}\leq\alpha\leq\alpha_{r^{*}}\right]\land\left[p_{s^{*}}\leq p^{\dagger}\right]\\ (p_{s},\left[p_{r^{*}},p^{\dagger}\right])&\left[\alpha_{s^{*},r^{*}}\leq\alpha\leq\alpha_{r^{*}}\right]\land\left[p^{\dagger}\leq p_{s^{*}}\right]\\ \end{cases}$
Case (ii), $c_{s}\geq c_{s^{*}}$:
(72) $(p_{r},p_{s})\in\begin{cases}(p_{r_{\text{ind}}},p_{r_{\text{ind}}})&\alpha\leq\alpha^{\dagger}_{r}\\ (p_{s},\left[p_{r^{*}},p_{s^{*}}\right])\cup([p_{s^{*}},p_{r_{\text{ind}}}],p_{s^{*}})&\left[\alpha^{\dagger}_{r}\leq\alpha\leq\alpha_{r^{*}}\right]\land\left[p_{s^{*}}\leq p^{\dagger}\right]\\ (p_{s},\left[p_{r^{*}},p^{\dagger}\right])&\left[\alpha^{\dagger}_{r}\leq\alpha\leq\alpha_{r^{*}}\right]\land\left[p^{\dagger}\leq p_{s^{*}}\right]\\ \end{cases}$
E.2. Relative Pareto optimality
Consider the equilibria in (70), (71), and (72), we will remove any that are Pareto suboptimal relative to other admissible Nash equilibria. There is one Pareto-suboptimal admissible equilibrium (indicated in the first case of (70) by Footnote 8), when $\alpha_{r^{*}}\leq\alpha\leq\alpha^{\dagger}_{s}$ and $p^{\dagger}\leq p_{s^{*}}$, in which case the retailer fulfilling demand at $p_{r^{*}}$ yields them the same payoff as the independent seller fulfilling demand at $p^{\dagger}$. However, the latter earns the independent seller positive payoff while the former earns the independent seller a payoff of 00. Thus, the equilibrium $(p_{r}=p_{r^{*}},p_{s}\in[p^{\dagger},p_{s^{*}}])$ is Pareto suboptimal relative to the shared-price equilibrium $p=p^{\dagger}$.
When $\alpha\leq\max\left\{\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\right\}$ or $\alpha\geq 1-c_{s}$, there is only one admissible equilibrium. When $\alpha^{\dagger}_{s}\leq\alpha\leq 1-c_{s}$, despite there being a continuum of equilibria in some of these cases, they all result in the same outcome with the retailer fulfilling demand at $p_{r^{*}}$. Similarly, when $\alpha_{s^{*}}\leq\alpha\leq\alpha_{s^{*},r^{*}}$, all of the equilibria result in the independent seller fulfilling demand at $p_{s^{*}}$. In the remaining cases, all equilibria are Pareto optimal, because the retailer prefers to push the price lower towards $p_{r^{*}}$ and/or $\frac{1}{2}$ (their optimal price when the independent seller fulfills demand) and the independent seller prefers to push the price higher towards $p_{s^{*}}$.
Appendix F Optimizing the referral fee - proofs
F.1. Nash equilibrium refinements
First, we justify the claims made in Section 5 about relative Pareto optimality and admissibility of equilibrium strategies for the sequential game.
Suppose there are two Nash equilibria of the sequential game $(\alpha_{1},p_{r},p_{s})$ and $(\alpha_{2},p_{r}^{\prime},p_{s}^{\prime})$ with $\alpha_{1}<\alpha_{2}$. Because these are both equilibria, they must have the same payoff to the retailer. However, because the independent seller’s payoff is monotonically decreasing in $\alpha$, $(\alpha_{2},p_{r}^{\prime},p_{s}^{\prime})$ is less Pareto optimal than $(\alpha_{1},p_{r},p_{s})$. Thus, Pareto optimality yields us the natural refinement that the retailer will choose the minimum $\alpha$ necessary to attain their maximum possible equilibrium payoff.
Let $\alpha_{1}<\alpha_{2}$, we will produce a strategy for the independent seller in which the retailer achieves a higher payoff with $\alpha_{1}$ and another strategy in which the retailer achieves a higher payoff with $\alpha_{2}$. Consider the strategy: “choose price $p_{s_{1}}$ if $\alpha\leq\frac{\alpha_{1}+\alpha_{2}}{2}$, otherwise choose price $p_{s_{2}}$” where $p_{s_{1}},p_{s_{2}}$ are chosen such that $\pi_{r,s}(p_{s_{1}},\alpha_{1})>\pi_{r,s}(p_{s_{2}},\alpha_{2})$. Switching the strategy to play $p_{s_{2}}$ when $\alpha\leq\frac{\alpha_{1}+\alpha_{2}}{2}$ and $p_{s_{1}}$ otherwise concludes the proof that both $\alpha_{1},\alpha_{2}$ are admissible.
F.2. Investigating equilibrium strategy profiles
Despite the preponderance of possible equilibrium strategy profiles, examining (26) and (27) shows there are only a few possible admissible, Pareto optimal outcomes in terms of who fulfills demand and what price they set. The only region of parameter space in which $\rho$ really matters is when $\alpha\in\left[\max\left\{\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\right\},\alpha^{\dagger}_{s}\right]$, because in this case there are a continuum of prices $p\in\left[\max\left\{p_{r^{*}},p_{s_{\text{ind}}}\right\},\min\left\{p_{s^{*}},p^{\dagger}\right\}\right]$ at which the independent seller could fulfill demand in equilibrium. Intuitively, this means that though $\rho$ could be very complicated in general, this is the only region in which this generality actually has an effect on the payoffs and the real-world outcomes.
We’ll denote as $\pi_{r}^{\text{(eq)}}(\rho(\alpha),c_{r},c_{s})$ the equilibrium payoff function for the retailer with strategy $\rho$ and $\pi_{s}^{\text{(eq)}}$ the analogous function for the independent seller. With this notation, we then have $\alpha_{*}\equiv\arg\max_{\alpha}\pi_{r}^{\text{(eq)}}(\rho(\alpha),c_{r},c_{s})$, and our eventual goal is to compute this $\alpha_{*}$. Even with $c_{r},c_{s},\rho$ held constant these are still complicated piecewise functions of $\alpha$, so as a first step we will prove continuity of the equilibrium payoffs in $\alpha$.
Proposition 0.
For all $\rho$ that are continuous on $\alpha\in\left(\max\left\{\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\right\},\alpha^{\dagger}_{s}\right)$, [^9] Throughout this paper we consider only pure strategies. However, consider for a moment mixed strategies $\mu_{\alpha}(p)$ such that for each $\alpha$, $\mu_{\alpha}$ is a valid density over equilibrium prices. Then, the analysis easily generalizes if we define $\mathbb{E_{\mu_{\alpha}}}[p]\equiv\rho(\alpha)$. $\pi_{r}^{\text{(eq)}}$ and $\pi_{s}^{\text{(eq)}}$ are continuous in $\alpha$.
Proof.
Examine the equilibrium prices in (26) and (27). Within each of the pieces of the function, $\pi_{r}^{\text{(eq)}},\pi_{s}^{\text{(eq)}}$ are constant as there is only one price at which demand will be fulfilled in equilibrium, except for when $\alpha\in\left(\max\left\{\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\right\},\alpha^{\dagger}_{s}\right)$. In this interval, we have
$\pi_{r}^{\text{(eq)}}(\alpha,c_{r},c_{s},\rho)=\pi_{r,s}(p_{s}=\rho(\alpha);c_{r},c_{s}),\qquad\pi_{s}^{\text{(eq)}}(\alpha,c_{r},c_{s},\rho)=\pi_{s,s}(p_{s}=\rho(\alpha);c_{r},c_{s})$
because $\rho$ is continuous in $\alpha$ on this interval by assumption and $\pi_{r,s},\pi_{s,s}$ are continuous functions of the price $p_{s}$, this is a composition of continuous functions which is continuous.
Thus, all that remains to check is that at each of the transitions between pieces, the lower limit is the same as the upper limit of $\pi_{r}^{\text{(eq)}}$ and $\pi_{s}^{\text{(eq)}}$. We showed in (50) that $p_{r_{\text{ind}}}=p_{s^{*}}$ when $\alpha=\alpha_{s^{*}}$. However, the remaining transitions are not quite as clear, as they depend on the value of $\rho$ when $\alpha=\max\left\{\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\right\}$ and $\alpha=\alpha^{\dagger}_{s}$. We will handle each of these cases in turn.
Now suppose that $c_{s}\geq c_{s^{*}}$. When $\alpha=\alpha^{\dagger}_{r}$, we showed in (63) that $p_{r^{*}}=p^{\dagger}$ (and $p_{r^{*}}<p^{\dagger}$ for $\alpha>\alpha^{\dagger}_{r}$). Because $\rho$ yields an admissible, Pareto optimal equilibrium price configuration for each $\alpha$ and $\pi_{r,s}$ is continous, the squeeze theorem tells us that
(73) $\pi_{r,s}(p_{r^{*}})=\lim_{\alpha\downarrow\alpha^{\dagger}_{r}}\pi_{r,s}(p_{r^{*}})\leq\lim_{\alpha\downarrow\alpha^{\dagger}_{r}}\pi_{r}^{\text{(eq)}}(\rho(\alpha))\leq\lim_{\alpha\downarrow\alpha^{\dagger}_{r}}\pi_{r,s}(p^{\dagger})=\pi_{r,s}(p_{r^{*}})$
Furthermore, we showed in (49) that at $\alpha=\alpha^{\dagger}_{r}$, $p_{r_{\text{ind}}}=p_{r^{*}}$, implying continuity because
(74) $\lim_{\alpha\uparrow\alpha^{\dagger}_{r}}\pi_{r}^{\text{(eq)}}(\rho(\alpha))=\lim_{\alpha\uparrow\alpha^{\dagger}_{r}}\pi_{r,s}(p_{r_{\text{ind}}})=\pi_{r,s}(p_{r_{\text{ind}}})=\pi_{r,s}(p_{r^{*}})=\lim_{\alpha\downarrow\alpha^{\dagger}_{r}}\pi_{r}^{\text{(eq)}}(\rho(\alpha))$
A similar application of the squeeze theorem shows continuity in when $c_{s}\leq c_{s^{*}}$ as $\alpha\to\alpha_{s^{*},r^{*}}$, because we showed in (52) that $p_{s^{*}}\leq p_{r^{*}}\iff\alpha\leq\alpha_{s^{*},r^{*}}$.
We know by construction that $\alpha^{\dagger}_{s}$ is the referral fee at which $p_{s_{\text{ind}}}=p^{\dagger}$. Because $p^{\dagger}$ was defined satisfy $\pi_{r,s}(p^{\dagger})=\pi_{r,r}(p_{r^{*}})$, and above $\alpha^{\dagger}_{s}$ the equlibrium outcome is the retailer fulfulling demand at $p_{r^{*}}$, we can apply the squeeze theorem a final time similar to show continuity as $\alpha\to\alpha^{\dagger}_{s}$. ∎
Corollary 0.
For $\alpha\in\left[\max\left\{\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\right\},\alpha^{\dagger}_{s}\right]$, let $\underline{\rho}(\alpha)\equiv\max\left\{p_{r^{*}},p_{s_{\text{ind}}}\right\}$ and $\overline{\rho}(\alpha)\equiv\min\left\{p_{s^{*}},p^{\dagger}\right\}$. Then, $\pi_{r}^{\text{(eq)}}(\overline{\rho}(\alpha))\leq\pi_{r}^{\text{(eq)}}(\rho(\alpha))\leq\pi_{r}^{\text{(eq)}}(\underline{\rho}(\alpha))$, and $\pi_{s}^{\text{(eq)}}(\underline{\rho}(\alpha))\leq\pi_{s}^{\text{(eq)}}(\rho(\alpha))\leq\pi_{s}^{\text{(eq)}}(\overline{\rho}(\alpha))$.
Proof.
Because $\rho$ is an equilibrium strategy profile, we know that for all $\alpha\in\left[\max\left\{\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\right\},\alpha^{\dagger}_{s}\right]$, $\rho(\alpha)\in\left[\max\left\{p_{r^{*}},p_{s_{\text{ind}}}\right\},\min\left\{p_{s^{*}},p^{\dagger}\right\}\right]\subset\left[\frac{1}{2},p_{s^{*}}\right]$. However, note that $\frac{1}{2}$ is the maximizer of the concave quadratic $\pi_{r,s}$ and $p_{s^{*}}$ is the maximizer of the concave quadratic $\pi_{s,s}$. Thus, for any given $\alpha\in\left[\max\left\{\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\right\},\alpha^{\dagger}_{s}\right]$, the retailer achieves their maximum (minimum) payoff by choosing the lowest (highest) possible equilibrium price and the independent seller achieves their maximum (minimum) payoff by choosing the highest (lowest) possible equilibrium price. ∎
F.3. Lower bounding the retailer’s payoff
Following the argument from Corollary 2, we know given $\alpha\in\left[\max\left\{\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\right\},\alpha^{\dagger}_{s}\right]$, the retailer’s payoff is minimized at the maximum possible price. In other words, the worst-case $\rho$ would be one that always chooses the maximum possible price when there are multiple equilibrium prices, precisely $\overline{\rho}(\alpha)\equiv\min\left\{p_{s^{*}},p^{\dagger}\right\}$ for $\alpha\in\left[\max\left\{\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\right\},\alpha^{\dagger}_{s}\right]$. Formally, for all $\rho$, we have
$\max_{\alpha}\pi_{r}^{\text{(eq)}}(\alpha,c_{r},c_{s},\rho)\geq\max_{\alpha}\pi_{r}^{\text{(eq)}}(\alpha,c_{r},c_{s},\overline{\rho})$
Thus, we turn our attention to maximizing the retailer’s equilibrium payoff in the worst case with $\rho=\overline{\rho}$. First, we show the maximizer of $\pi_{r}^{\text{(eq)}}$ is never in a piece where the independent seller fulfills demand at $p_{r_{\text{ind}}}$, because the retailer’s equilibrium payoff is never greater than their payoff when setting the fee above $\overline{\alpha}$ and fulfilling demand themselves.
Lemma 0.
$\pi_{r,s}(p_{r_{\text{ind}}},\alpha)\leq\pi_{r,r}(p_{r^{*}})$
Proof.
First, let’s find the derivative of $\pi_{r,s}(p_{r_{\text{ind}}},\alpha)$:
$\frac{\partial\pi_{r,s}(p_{r_{\text{ind}}},\alpha)}{\partial\alpha}=\frac{\partial}{\partial\alpha}\left(\frac{\alpha c_{r}(1-\alpha-c_{r})}{(1-\alpha)^{2}}\right)=\frac{c_{r}[1-c_{r}-\alpha(1+c_{r})]}{(1-\alpha)^{3}}$
Notice that $\frac{\partial\pi_{r,s}(p_{r_{\text{ind}}},\alpha)}{\partial\alpha}\geq 0$ for all $\alpha\leq\min\left\{\alpha_{s^{*}},\alpha^{\dagger}_{r}\right\}$, and that there is a maximum exactly when $\alpha=\frac{1-c_{r}}{1+c_{r}}=\alpha^{\dagger}_{r}$. Substituting this value into $\pi_{r,s}(p_{r_{\text{ind}}},\alpha)$, we find that
$\pi_{r,s}(p_{r_{\text{ind}}},\alpha)$ $\leq\pi_{r,s}(p_{r_{\text{ind}}},\alpha^{\dagger}_{r})$
$=\frac{\left(\frac{1-c_{r}}{1+c_{r}}\right)c_{r}\left(\left(\frac{2c_{r}}{1+c_{r}}\right)-c_{r}\right)}{\left(\frac{2c_{r}}{1+c_{r}}\right)^{2}}$
$=\frac{c_{r}\left(\frac{1-c_{r}}{1+c_{r}}\right)\left(\frac{c_{r}-c_{r}^{2}}{1+c_{r}}\right)}{\left(\frac{2c_{r}}{1+c_{r}}\right)^{2}}$
$=\left(\frac{1-c_{r}}{2}\right)^{2}$
$=\pi_{r,r}(p_{r^{*}})$
Because we are working with $\overline{\rho}$, the only remaining equilibrium outcomes that could maximize the retailer’s payoff are ones where the independent seller fulfills demand at $p_{s^{*}}$ or $p^{\dagger}$ and one where the retailer fulfills demand at $p_{r^{*}}$. However, by construction $\pi_{r,s}(p^{\dagger})=\pi_{r,r}(p_{r^{*}})$. Thus, all that remains is to understand if or when the retailer can achieve a greater payoff than $\pi_{r,r}(p_{r^{*}})$ in an equilibrium where the independent seller fulfills demand at $p_{s^{*}}$.
To this end, we will first find the $\alpha$ that maximizes $\pi_{r,s}(p_{s^{*}},\alpha)$, temporarily ignoring the question of if that $\alpha$ is actually an equilibrium where $p_{s^{*}}$ is observed, which will be addressed separately in Lemma 5.
Lemma 0.
[…]
Proof.
We found the derivative of $\pi_{r,s}(p_{s^{*}},\alpha)$ in (64). Setting the expression equal to 00, we find that the critical value $\overline{\alpha}$ must satisfy
(75) $c_{s}^{2}=\frac{(1-\overline{\alpha})^{3}}{1+\overline{\alpha}}$
We will explicitly solve the cubic (75) for $\overline{\alpha}$. As an important remark that is worth restating, the function $f(x)=\frac{(1-x)^{3}}{1+x}$ is a bijection from $(0,1)\to(0,1)$, so $\overline{\alpha}$ is the unique value that satisfies (64) for $c_{s}^{2}\in(0,1)$. To obtain our explicit solution, we first write it in standard form:
$c_{s}^{2}$ $=\frac{(1-\overline{\alpha})^{3}}{1+\overline{\alpha}}$
$c_{s}^{2}+c_{s}^{2}\overline{\alpha}$ $=-\overline{\alpha}^{3}+3\overline{\alpha}^{2}-3\overline{\alpha}+1$
00 0 $=\overline{\alpha}^{3}-3\overline{\alpha}^{2}+(3+c_{s}^{2})\overline{\alpha}+c_{s}^{2}-1$
We perform the change of variables $t\equiv\overline{\alpha}-1$ to obtain a depressed cubic:
$t^{3}+c_{s}^{2}t+2c_{s}^{2}=0\iff 0=\overline{\alpha}^{3}-3\overline{\alpha}^{2}+(3+c_{s}^{2})\overline{\alpha}+c_{s}^{2}-1$
Using Cardano’s formula to solve the depressed cubic, we find that the real root is
[…]
as desired. Note that the discriminant is positive, so $\overline{\alpha}$ is indeed always real. ∎
Note that if there exists any $\alpha$ such that $p_{s^{*}}\leq p^{\dagger}$, then $\overline{\alpha}$ also satisfies $p_{s^{*}}\leq p^{\dagger}$. This follows because $\pi_{r,s}$ is a concave quadratic maximized at $p=\frac{1}{2}$ for all $\alpha$, and because $\frac{1}{2}\leq p_{s^{*}},p^{\dagger}$, we have
(76) $p_{s^{*}}\leq p^{\dagger}\iff\pi_{r,s}(p^{\dagger})\leq\pi_{r,s}(p_{s^{*}},\alpha)\leq\pi_{r,s}(p_{s^{*}},\overline{\alpha})$
Before proceeding, we’ll pause to build some intuition about when this condition holds.
Remark 1.
If $c_{s}\leq c_{s^{*}}$, $\pi_{r,s}(p_{s^{*}},\overline{\alpha})\geq\pi_{r,s}(p^{\dagger})$ always.
Proof.
First, note that
$\pi_{r,s}(p^{\dagger})$ $\leq\pi_{r,s}(p_{s^{*}},\alpha)$
$\left(\frac{1-c_{r}}{2}\right)^{2}$ $\leq\frac{\alpha}{4}\left(1-\left(\frac{c_{s}}{1-\alpha}\right)^{2}\right)$
(77) $c_{r}$ $\geq 1-\sqrt{\alpha\left(1-\left(\frac{c_{s}}{1-\alpha}\right)^{2}\right)}$
Consider $\alpha=\alpha_{s^{*},r^{*}}$, which we showed in (53) and (61) satisfies $\alpha_{s^{*},r^{*}}\leq\alpha_{r^{*}}\leq\alpha^{\dagger}_{s}$. We will show that the condition (77) always holds when $c_{s}\leq c_{s^{*}}$. Indeed, recalling from (54) that $c_{s}\leq c_{s^{*}}$ is equivalent to $\alpha_{s^{*},r^{*}}\geq\alpha^{\dagger}_{r}$, we have
$1-\sqrt{\alpha_{s^{*},r^{*}}\left(1-\left(\frac{c_{s}}{1-\alpha_{s^{*},r^{*}}}\right)^{2}\right)}$ $=1-\sqrt{\left(\frac{c_{r}-c_{s}}{c_{r}}\right)\left(1-\left(\frac{c_{s}}{\frac{c_{s}}{c_{r}}}\right)^{2}\right)}$
$=1-\sqrt{\left(\frac{c_{r}-c_{s}}{c_{r}}\right)\left(1-c_{r}^{2}\right)}$
$\leq 1-\sqrt{\left(\frac{1-c_{r}}{1+c_{r}}\right)(1-c_{r})(1+c_{r})}$
$=1-|1-c_{r}|=c_{r}$
The proof of Remark 1 breaks down when $c_{s}\geq c_{s^{*}}$ because we have now that $\alpha_{s^{*},r^{*}}\leq\alpha^{\dagger}_{r}$. When $c_{s}\geq c_{s^{*}}$, there sometimes exists an $\alpha$ satisfying (76) and sometimes does not. With this, we can now answer the question of if this maximum payoff is actually attainable in equilibrium.
Lemma 0.
If $\pi_{r,s}(p_{s^{*}},\overline{\alpha})\geq\pi_{r,s}(p^{\dagger})$, then $\overline{\alpha}\in[\alpha_{s^{*}},\alpha^{\dagger}_{s}]$.
Proof.
To prove the upper bound, suppose for contradiction that $\overline{\alpha}>\alpha^{\dagger}_{s}$. Then, by construction of $\alpha^{\dagger}_{s}$ and (76), we have a contradiction because
$p^{\dagger}<p_{s_{\text{ind}}}<p_{s^{*}}=\frac{1+p_{s_{\text{ind}}}}{2}\iff\pi_{r,s}(p^{\dagger})>\pi_{r,s}(p_{s^{*}})$
For the lower bound, we prove a stronger statement. First, note that for all $\alpha\in(0,1]$,
(78) $\alpha\geq\alpha_{s^{*}}=1-2c_{r}+c_{s}\iff c_{r}\geq\frac{1+c_{s}-\alpha}{2}=1-\frac{1+\alpha-c_{s}}{2}$
We claim that, for all $\alpha\in(0,1]$,
(79) $1-\sqrt{\alpha\left(1-\left(\frac{c_{s}}{1-\alpha}\right)^{2}\right)}\geq 1-\frac{1+\alpha-c_{s}}{2}$
Combining (79) with (76), (77), (78) is sufficient to conclude the proof because
$\pi_{r,s}(p_{s^{*}},\overline{\alpha})\geq\pi_{r,s}(p^{\dagger})\implies c_{r}\geq 1-\sqrt{\overline{\alpha}\left(1-\left(\frac{c_{s}}{1-\overline{\alpha}}\right)^{2}\right)}\implies c_{r}\geq\frac{1-\overline{\alpha}+c_{s}}{2}\implies\overline{\alpha}\geq 1-2c_{r}+c_{s}$
Thus, all that remains is to show (79). Manipulating the inequality, we have
$1-\sqrt{\alpha\left(1-\left(\frac{c_{s}}{1-\alpha}\right)^{2}\right)}$ $\geq 1-\frac{1+\alpha-c_{s}}{2}$
(80) $\frac{1-c_{s}+\alpha}{2}$ $\geq\sqrt{\alpha\left(1-\left(\frac{c_{s}}{1-\alpha}\right)^{2}\right)}$
$(1-c_{s})^{2}+\alpha^{2}+2\alpha(1-c_{s})$ $\geq 4\alpha-\frac{4\alpha c_{s}^{2}}{(1-\alpha)^{2}}$
$(\alpha^{2}-2\alpha+1)\left[(1-c_{s})^{2}+\alpha^{2}+2\alpha(1-c_{s})\right]$ $\geq 4\alpha(\alpha^{2}-2\alpha+1)-4\alpha c_{s}^{2}$
(81) $\alpha^{4}-2(2+c_{s})\alpha^{3}+(c_{s}^{2}+2c_{s}+6)\alpha^{2}-2(c_{s}+2)(1-c_{s})\alpha+(1-c_{s})^{2}$ $\geq 0$
To minimize the left side, we first take the derivative with respect to $c_{s}$ to find the critical $\tilde{c}_{s}$
$\frac{\partial}{\partial c_{s}}$ $\left[\alpha^{4}-2(2+c_{s})\alpha^{3}+(c_{s}^{2}+2c_{s}+6)\alpha^{2}-2(c_{s}+2)(1-c_{s})\alpha+(1-c_{s})^{2}\right]$
$=2c_{s}(1+\alpha)^{2}+2(-\alpha^{3}+\alpha^{2}+\alpha-1)\stackrel{{\scriptstyle\text{set}}}{{=}}0$
$\iff\tilde{c}_{s}$ $=\frac{\alpha^{3}+1-\alpha(1+\alpha)}{1+\alpha}=\frac{(\alpha^{2}-2\alpha+1)(1+\alpha)-\alpha(1+\alpha)}{1+\alpha}=\frac{(1-\alpha)^{2}}{1+\alpha}$
Thus, the left side of (81) is lower bounded by its value at $\tilde{c}_{s}=\frac{(1-\alpha)^{2}}{1+\alpha}$. We will show that (80) holds with equality for $\tilde{c}_{s}$, which is sufficient to conclude the proof. Indeed, for the left side we have
$\frac{1+\alpha-\tilde{c}_{s}}{2}$ $=\frac{1+\alpha-\frac{(1-\alpha)^{2}}{1+\alpha}}{2}$
$=\frac{(1+\alpha)^{2}-(1-\alpha)^{2}}{2(1+\alpha)}$
$=\frac{2\alpha}{1+\alpha}$
and for the right side we have
$\sqrt{\alpha\left(1-\left(\frac{\tilde{c}_{s}}{1-\alpha}\right)^{2}\right)}$ $=\sqrt{\alpha\left(1-\left(\frac{\frac{(1-\alpha)^{2}}{1+\alpha}}{1-\alpha}\right)^{2}\right)}$
$=\sqrt{\alpha\left(1-\left(\frac{1-\alpha}{1+\alpha}\right)^{2}\right)}$
$=\frac{2\alpha}{1+\alpha}$
With this, we have shown that if the retailer chooses $\overline{\alpha}$, $\overline{\rho}$ yields an equilibrium strategy profile in which the independent seller fulfills demand at price $p_{s^{*}}$. This is sufficient to conclude that:
(82) $\max_{\alpha}\pi_{r}^{\text{(eq)}}(\alpha,c_{r},c_{s},\rho)\geq\max_{\alpha}\pi_{r}^{\text{(eq)}}(\alpha,c_{r},c_{s},\overline{\rho})=\begin{cases}\pi_{r,s}(p_{s^{*}},\overline{\alpha})&\exists\alpha\text{ s.t. }\pi_{r,s}(p_{s^{*}},\alpha)>\pi_{r,r}(p_{r^{*}})\\ \pi_{r,r}(p_{r^{*}})&\text{else}\end{cases}$
We have shown that $\alpha_{*}(\overline{\rho})=\overline{\alpha}$ in the first case and $\alpha_{*}(\overline{\rho})=\max\left\{\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\right\}$ otherwise. Ideally, we would be able to directly translate this to a tight bound on $\alpha_{*}\geq\overline{\alpha}$ for all $\rho$ in the first case, however it is still possible that there exists some $\rho$ and some $\alpha^{\prime}<\overline{\alpha}$ such that
(83) $\max_{\alpha}\pi_{r,s}(\rho(\alpha))=\pi_{r,s}(\rho(\alpha^{\prime}))>\pi_{r,s}(\rho(\overline{\alpha}))$
For instance, (83) could hold if $\rho(\alpha^{\prime})=p_{r^{*}}$ for some well-chosen $\alpha^{\prime}<\overline{\alpha}$, but for all other $\alpha$, we set $\rho=\overline{\rho}$. If we impose some assumptions on the functional form of $\rho$, we may be able to translate (82) into a tighter bound on $\alpha_{*}$. However, we do not investigate further tightening the bound in this work as we have at least shown that the lower bound $\alpha_{*}\geq\max\left\{\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\right\}$ is tight whenever we are in the second case of (82).
With this, the maximum payoff that the independent seller could achieve in equilibrium of the sequential game is either by fulfilling demand at $p^{\dagger}$ or $p_{s^{*}}$. If it is at $p_{s^{*}}$, clearly they would prefer the lowest possible $\alpha$. If it is at $p^{\dagger}$ for some $\alpha^{\prime}\geq\max\left\{\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\right\}$, we have
$\pi_{s,s}\left(p^{\dagger},\alpha^{\prime}\right)\leq\pi_{s,s}\left(p^{\dagger},\max\left\{\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\right\}\right)\leq\pi_{s,s}\left(p_{s^{*}},\max\left\{\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\right\}\right)$
This leaves us with the result that the independent seller’s payoff in equilibrium of the sequential game is bounded by
$\pi_{s}^{\text{(eq)}}(\rho(\alpha_{*}))\leq\pi_{s,s}\left(p_{s^{*}},\max\left\{\alpha^{\dagger}_{r},\alpha_{s^{*},r^{*}}\right\}\right)$
which is loose for similar reasons as our bound on $\alpha_{*}$.
F.4. Upper bounding the retailer’s payoff
Analogous to the previous subsection, we know that the upper bound for the retailer’s payoff will be with the strategy profile $\underline{\rho}\equiv\max\left\{p_{r^{*}},p_{s_{\text{ind}}}\right\}$. The retailer’s maximum payoff either occurs when the independent seller fulfills demand at $p_{r^{*}}$ or $p_{s_{\text{ind}}}$, or when the retailer fulfills demand themselves at $p_{r^{*}}$. Recall that $p_{r^{*}}\leq p_{s_{\text{ind}}}\iff\alpha\leq\alpha_{r^{*}}$ per (46). The retailer’s payoff function when the independent seller fulfills at $p_{r^{*}}$ is simple –
$\pi_{r,s}(p_{r^{*}},\alpha)=\alpha p_{r^{*}}(1-p_{r^{*}})=\frac{\alpha}{4}(1-c_{r}^{2})$
This is clearly increasing in $\alpha$, so the maximum payoff that the retailer could achieve when the independent seller fulfills at $p_{r^{*}}$ is by setting $\alpha=\alpha_{r^{*}}$. At this $\alpha$, we have
(84) $\pi_{r,s}(p_{r^{*}},\alpha_{r^{*}})=\frac{1-2c_{s}+c_{r}}{4(1+c_{r})}\left(\frac{1-c_{r}}{2}\right)^{2}=\frac{1+c_{r}-2c_{s}}{4(1+c_{r})}\left(1-c_{r}^{2}\right)=\left(\frac{1+c_{r}-2c_{s}}{2}\right)\left(\frac{1-c_{r}}{2}\right)$
Notice that this yields a higher payoff than the retailer fulfilling demand on their own, because
$\left(\frac{1+c_{r}-2c_{s}}{2}\right)\left(\frac{1-c_{r}}{2}\right)$ $>\left(\frac{1-c_{r}}{2}\right)^{2}=\pi_{r,r}(p_{r^{*}})$
$1+c_{r}-2c_{s}$ $>1-c_{r}$
$c_{r}$ $>c_{s}$
Finally, it remains to see if the retailer could achieve a higher payoff than (84) for $\alpha\in\left[\alpha_{r^{*}},\alpha^{\dagger}_{s}\right]$ at price $p_{s_{\text{ind}}}$. The payoff function is slightly less simple –
$\pi_{r,s}(p_{s_{\text{ind}}},\alpha)=\alpha\left(\frac{c_{s}}{1-\alpha}\right)\left(1-\frac{c_{s}}{1-\alpha}\right)$
it’s not obvious if this is increasing or decreasing in $\alpha$, so we will take the derivative:
(85) $\frac{\partial\pi_{r,s}(p_{s_{\text{ind}}},\alpha)}{\partial\alpha}=\frac{\partial}{\partial\alpha}\left(\frac{\alpha c_{s}(1-\alpha-c_{s})}{(1-\alpha)^{2}}\right)=\frac{c_{s}[1-c_{s}-\alpha(1+c_{s})]}{(1-\alpha)^{3}}$
Clearly, the unique maximum occurs when $\alpha=\frac{1-c_{s}}{1+c_{s}}$, and $\pi_{r,s}(p_{s_{\text{ind}}},\alpha)$ is decreasing in $\alpha$ for $\alpha>\frac{1-c_{s}}{1+c_{s}}$. However, we have that
(86) $\frac{1-c_{s}}{1+c_{s}}$ $<1-\frac{2c_{s}}{1+c_{r}}=\alpha_{r^{*}}$
$\frac{2c_{s}}{1+c_{r}}$ $<\frac{2c_{s}}{1+c_{s}}$
$c_{s}$ $<c_{r}$
Thus, the maximum equilibrium payoff the retailer can attain when the retailer sells at $p_{s_{\text{ind}}}$ is at the minimum possible $\alpha$, which is in fact $\alpha=\alpha_{r^{*}}$. Conventiently, this implies that the maximizer of the retailer’s payoff function with the strategy $\underline{\rho}$ is precisely $\underline{\alpha}\equiv\alpha_{r^{*}}$. Because at $\underline{\alpha}$, we have $p_{s_{\text{ind}}}=p_{r^{*}}$, this is sufficient to conclude that
$\max_{\alpha}\pi_{r}^{\text{(eq)}}(\alpha,c_{r},c_{s},\rho)\leq\max_{\alpha}\pi_{r}^{\text{(eq)}}(\alpha,c_{r},c_{s},\underline{\rho})=\pi_{r}^{\text{(eq)}}(\underline{\alpha},c_{r},c_{s},\underline{\rho})=\left(\frac{1+c_{r}-2c_{s}}{2}\right)\left(\frac{1-c_{r}}{2}\right)$
Just as the previous case, this is unfortunately not sufficient to conclude that $\alpha_{*}\leq\underline{\alpha}$ in general; the best we can do is conclude that $\alpha_{*}\leq\alpha^{\dagger}_{s}$. For the independent seller, it proves the trivial bound that $\pi_{s}^{\text{(eq)}}(\alpha,c_{r},c_{s},\rho)\geq 0$, where the bound is tight because when $\alpha=\underline{\alpha}$ and $\rho=\underline{\rho}$, the independent seller is fulfilling demand at $p_{s_{\text{ind}}}$ and achieves a payoff of 00.
Appendix G Equilibria of leaving subgame
The leaving subgame is simply a standard Bertrand game with asymmetric costs $c_{s}+\delta<c_{r}$. Though this has been derived in the literature for general demand curves such as in (Blume, 2003; Kartik, 2011), for completeness we include here a brief derivation of the equilibria contextualized to our setting. First, note that $\beta\in\{0,1\}$, which follows as a corollary of Proposition 1 when $\alpha=0$. Thus, any shared-price equilibrium must have exactly one player fulfill demand. However, we have that
Lemma 0.
In all Nash equilibria of the leaving subgame, the independent seller fulfills demand.
Proof.
Suppose for contradiction there exists such an equilibrium $(p_{r},p_{s}\geq p_{r})$. Clearly, we must have $p_{r}\geq c_{r}$, otherwise at a price $p_{r}^{\prime}<c_{r}$ the retailer would prefer to increase their price to $p_{s}+\varepsilon$ and achieve a payoff of 00 rather than achieving negative payoff for each unit sold at $p_{r}^{\prime}$. However, this implies that $p_{s}\geq c_{r}>c_{s}+\delta$, so setting any price $p_{s}^{\prime}\in(c_{s}+\delta,c_{r})$ yields the independent seller a payoff of
$0=\pi_{s}^{(\ell)}(p_{r},p_{s})<\pi_{s}^{(\ell)}(p_{r},p_{s}^{\prime})=(p_{s}^{\prime}-(c_{s}+\delta))(1-p_{s}^{\prime})$
With this, consider an equilibrium $(p_{r}\geq p_{s},p_{s})$ in which the independent seller fulfills demand. In order for this to be an equilibrium, we must have that
$p_{s}=\arg\max_{p_{s}^{\prime}\in[0,p_{r}]}\pi_{s}^{(\ell)}(p_{r},p_{s}^{\prime})$
However, because $\pi_{s}^{(\ell)}$ a concave quadratic maximized at $p_{s^{*}}^{(\ell)}\equiv\frac{1+(c_{s}+\delta)}{2}$, it follows that the equilibrium $p_{s}$ is given by
$p_{s}=\begin{cases}p_{r}&p_{r}\leq p_{s^{*}}^{(\ell)}\\ p_{s^{*}}^{(\ell)}&p_{s^{*}}^{(\ell)}\leq p_{r}\end{cases}$
Consider the first case where $p_{r}\leq p_{s^{*}}^{(\ell)}$. If $p_{r}>p_{s}$, because $\pi_{s}^{(\ell)}$ is increasing for $p_{s}\in(0,p_{s^{*}})\supset(0,p_{r}]$, the independent seller achieves a higher payoff by setting price $p_{s}+\varepsilon$ for some $\varepsilon\in(0,p_{r}-p_{s})$. We therefore must have a shared price $p_{r}=p=p_{s}$ in this case. If $p>c_{r}$, the retailer achieves a higher price by setting $p^{\prime}\in(c_{r},p)$ and fulfilling some demand, yielding them a nonzero payoff. Furthermore, if $p>p_{s^{*}}^{(\ell)}$, the independent seller would prefer to fulfill at their optimal price $p_{s^{*}}^{(\ell)}$ instead. Thus, we have shared-price Nash equilibria at $p\in\left[c_{s}+\delta,\min\left\{p_{s^{*}}^{(\ell)},c_{r}\right\}\right]$.
Now consider the second case where $p_{s^{*}}^{(\ell)}\leq p_{r}$. If $c_{r}\leq p_{s^{*}}^{(\ell)}$, our analysis reduces to that of the previous case. If, on the other hand, $c_{r}>p_{s^{*}}^{(\ell)}$, then we still have similar shared-price equilibria $(p\in[c_{s}+\delta,p_{s^{*}}^{(\ell)}],s)$. However, we also gain unequal-price Nash equilibria $(p_{r}>p_{s},p_{s}=p_{s^{*}}^{(\ell)})$ – the independent seller would no longer benefit from deviating to a price $p_{s}+\varepsilon$ and the retailer still would not prefer to undercut the independent seller because $p_{s^{*}}^{(\ell)}<c_{r}$.
To summarize, we have found the following Nash equilibria:
$\begin{cases}p\in\left[c_{s}+\delta,c_{r}\right]&c_{r}\leq p_{s^{*}}^{(\ell)}\\ \left(p\in\left[c_{s}+\delta,p_{s^{*}}^{(\ell)}\right]\right)\cup\left(p_{r}>p_{s},p_{s}=p_{s^{*}}^{(\ell)}\right)&p_{s^{*}}^{(\ell)}\leq c_{r}\end{cases}$
Now that we have found the Nash equilibria, we apply the admissibility refinement that we have been using throughout. Very similar analysis to Appendix E.1 shows that the admissible strategy sets for the retailer and seller are $p_{r}\in[c_{r},p_{r^{*}}]$ and $p_{s}\in\left[c_{s}+\delta,p_{s^{*}}^{(\ell)}\right]$. While the independent seller is not playing any dominated strategies in equilibrium, the retailer is, so this is a nontrivial refinement. The admissible Nash equilibria are:
$(p_{r},p_{s})=\begin{cases}c_{r}&c_{r}\leq p_{s^{*}}^{(\ell)}\\ \left(p_{r}\in[c_{r},p_{r^{*}}],p_{s}=p_{s^{*}}^{(\ell)}\right)&p_{s^{*}}^{(\ell)}\leq c_{r}\end{cases}$
From here, we find that the admissible Nash equilibrium outcomes are:
$(p,x)=\begin{cases}(c_{r},s)&c_{r}\leq p_{s^{*}}^{(\ell)}\\ (p_{s^{*}}^{(\ell)},s)&p_{s^{*}}^{(\ell)}\leq c_{r}\end{cases}$
or simply $(p,x)=\left(\min\left\{c_{r},p_{s^{*}}^{(\ell)}\right\},s\right)$.